### v1

# The Formality of the Goldman–Turaev Lie Bialgebra on a Closed Surface

Abstract
We reformulate the Kashiwara–Vergne groups and associators in higher genera, introduced in Alekseev–Kawazumi–Kuno–Naef, in terms of non-commutative connections using the tools developed in a previous paper. As the main result, the case of closed surfaces is dealt with to determine the pro-nilpotent automorphism group of the associated graded of the Goldman–Turaev Lie bialgebra.
2020 Mathematics Subject Classification: 16D20, 17A61, 53D30, 57K20, 58B34.
Keywords: loop operations, the Goldman–Turaev Lie bialgebra, formality problem, non-commutative geometry, divergence maps, flat connections.
1. Introduction
The Kashiwara–Vergne (KV) problem originated from Lie theory asking the existence of certain infinite Lie series, the Kashiwara–Vergne associators. The problem is partially solved in their original paper [KV78] and then by Rouvière [Rou81] and Vergne [Ver99], and completely solved by Alekseev–Meinrenken [AM06]. Later, by a remarkable result by Alekseev–Torossian [AT12], the KV problem is reduced to the existence of Drinfeld associators, which is known over the field of characteristic zero in the original paper [Dri90] by Drinfeld. The space of solutions to the KV problem is a bi-torsor over the groups KV and KRV, which are called the Kashiwara–Vergne groups.
The defining relations of the KV associators and the KV groups have a higher genus analogue. Alekseev–Kawazumi–Kuno–Naef [AKKN23] shows that, in the genus 00 0 case, the formality problem of the Goldman–Turaev Lie bialgebra is essentially equivalent to the KV problem; this perspective immediately allows us to define the higher genus KV associators as the solutions to the formality problem and the KV groups as the (pro-unipotent) automorphism group of the (graded) Goldman–Turaev Lie bialgebra. On the other hand, we have higher genus analogues of the Drinfeld associators. For genus 11 11, they are called elliptic associators introduced by Enriquez [Enr14], and their relation to the corresponding KV problem is discussed in [AKKN23]. For an arbitrary genus, several generalisations of the GT groups and the Drinfeld associators are proposed: [Gon20] by Gonzalez, [Fel21] by Felder, and Section 5 of [CIW19] by Campos–Idrissi–Willwacher. They all agree in genus 11 11, but the relation between them and the KV associators for genus $\geq 2$ is still an open question.
Getting back to the KV problem in higher genera, the existence problem itself is completely solved in [AKKN23], and the key step in the proof is the factorisation of the Turaev cobracket in two parts: a Hamiltonian flow and a non-commutative divergence. In this paper, we solve the formality problem on a closed surface using the factorisation of the Turaev cobracket obtained in the author’s previous paper [Tan24b]. We start with a reformulation of the KV groups and the space of solution $\mathrm{SolKV}$ for a connected oriented compact surface $\Sigma$ with non-empty boundary in terms of connections in non-commutative geometry and later see the case of closed surfaces based on the observation.
Let $\mathbb{K}$ be a field of characteristic zero, $\pi=\pi_{1}(\Sigma)$ the fundamental group of $\Sigma$ and $H=H_{1}(\Sigma;\mathbb{K})$ the first homology group. In their formulation, the KV groups are described in terms of group 11 11-cocycles on the group $\operatorname{tAut}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\pi})$ of tangential automorphisms on the complete Hopf algebra $\widehat{\mathbb{K}\pi}$, and the graded counterpart $\operatorname{tAut}_{\mathrm{Hopf}}(\hat{T}(H))$ on the completed tensor algebra $\hat{T}(H)$. We first see that a tangential automorphism of $\widehat{\mathbb{K}\pi}$ is interpreted as a derivation of the complete Hopf groupoid (in the sense of Fresse [Fre17]) $\widehat{\mathbb{K}\mathscr{G}}$ associated with the fundamental groupoid $\mathscr{G}$ of $\Sigma$, and similarly for the graded counterpart $\hat{\mathscr{T}}$. We also have their subcategories $\partial\widehat{\mathbb{K}\mathscr{G}}$ and $\partial\hat{\mathscr{T}}$ generated by boundary loops.
Next, we define connections $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ and $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on some $\mathbb{K}\mathscr{G}$-module and $\hat{\mathscr{T}}$-module, respectively, depending on a framing $\mathsf{fr}$ on $\Sigma$ and a free-generating system $\mathcal{C}$ of $\pi$. Recall that a connection on a $A$-module $N$ is a linear map $\nabla\colon N\to\Omega^{1}A\otimes_{A}N$ satisfying the Leibniz rule, where $\Omega^{1}A$ is the space of non-commutative 11 11-forms. Then, the value of those 11 11-cocycles on an automorphism $G$ is written as (the trace of) the difference of connections by the action of $G$: $\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla-\nabla)$ where $\nabla$ is either of the two above. This takes its value in the space of de Rham 1-forms $\operatorname{DR}^{1}\!A$, a quotient space of $\Omega^{1}A$. We have the exterior derivative $d$, which sends an element of the cyclic quotient $|A|=A/[A,A]$ to $\operatorname{DR}^{1}\!A$. One of the main results is the following.
Theorem (Theorem 4.21).
The KV groups and the set of KV associators are expressed as follows:
  • $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})\in d|\partial\widehat{\mathbb{K}\mathscr{G}}|\Big{\}}$,
  • $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!\!_{H,\mathsf{fr}})\in d|\partial\hat{\mathscr{T}}|\Big{\}}$, and
  • $\mathrm{SolKV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\mathrm{Isom}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}},\widehat{\mathbb{K}\mathscr{G}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})\in d|\partial\widehat{\mathbb{K}\mathscr{G}}|\Big{\}}$.
Here, the subscript + indicates the pro-unipotent part of each space. In particular, the Lie algebra corresponding to $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$ is given by
$\mathfrak{krv}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}g\in\operatorname{Der}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}}):\operatorname{div}^{\nabla^{\prime}\!\!_{H,\mathsf{fr}}}(g)\in|\partial\hat{\mathscr{T}}|\Big{\}}.$
The divergence map $\operatorname{div}$ is defined from a connection, whose construction is given in [Tan24b].
Now consider the case of closed surfaces. The existence of a formality morphism follows from the case of a surface with only one boundary component, which is solved in [AKKN23], so it is enough to consider the automorphism group of the associated graded. Based on the reformulation above, we have a similar description in the case of closed surfaces:
Theorem (Theorem 6.2).
The pro-unipotent part of the automorphism group of the associated graded of the Goldman–Turaev Lie bialgebra is given by $\mathrm{KRV}_{(g,0)}=\operatorname{exp}(\mathfrak{krv}_{(g,0)})$, where
$\mathfrak{krv}_{(g,0)}:=\{g\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})\colon\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)\in\operatorname{Ker}(|\bar{\Delta}_{\omega}|)\}\,.$
Here, $\hat{L}(H)_{\omega}$ is some quotient of the completed free Lie algebra over $H$, $\nabla^{\prime}_{\bullet,H}$ is a connection on a $\hat{L}(H)_{\omega}$-module and $\bar{\Delta}_{\omega}\colon\hat{T}(H)_{\omega}\to\hat{T}(H)_{\omega}\hat{\otimes}\hat{T}(H)_{\omega}$ is the reduced coproduct: $\bar{\Delta}_{\omega}(x)=\Delta_{\omega}(x)-x\otimes 1-1\otimes x$.
The space $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)$ deserves some description, so we computed its lower degree parts at the end of the paper.
Organisation of the paper. In Sections 2 and 3, we recall the definition of Goldman–Turaev Lie bialgebra, the formality problem, and main results in [AKKN23]. Section 4 is occupied with the reformulation of the KV groups. In Section 5, we construct a basis of $|T(H)_{\omega}|$ for later sections. The formality problem on a closed surface is dealt with in Section 6. Section 7 is devoted to the computation of the kernel of the reduced coproducts.
Acknowledgements. The author thanks Yusuke Kuno for a discussion on Section 7 and Nariya Kawazumi for thoroughly reading a draft.
Conventions. $\mathbb{K}$ is a field of characteristic zero. Unadorned tensor products are always over $\mathbb{K}$.
2. The Goldman–Turaev Lie Bialgebra
In this section, we recall the Goldman–Turaev Lie bialgebra and its associated graded following [AKKN23]. Let $g\geq 0$, $n\geq-1$ and $\Sigma=\Sigma_{g,n+1}$ a connected compact oriented surface of genus $g$ and $n+1$ boundary components $\partial_{0}\Sigma,\dotsc,\partial_{n}\Sigma$. Consider the fundamental group $\pi=\pi_{1}(\Sigma,*)$ of $\Sigma$ and the group algebra $\mathbb{K}\pi$. We require the base point * to be taken on $\partial_{0}\Sigma$ if $n\geq 0$. For a $\mathbb{K}$-algebra $A$, we denote the cyclic quotient $A/[A,A]$ by $|A|$. Then, $|\mathbb{K}\pi|=\mathbb{K}\pi/[\mathbb{K}\pi,\mathbb{K}\pi]$ is identified with the $\mathbb{K}$-vector space spanned by the homotopy set of free loops on $\Sigma$.
Definition 2.1.
We define the Goldman bracket $[\cdot,\cdot]\colon|\mathbb{K}\pi|^{\otimes 2}\to|\mathbb{K}\pi|$ by, for generically immersed free loops $\alpha$ and $\beta$ on $\Sigma$,
$[\alpha,\beta]=\sum_{p\in\alpha\cap\beta}\mathrm{sign}(p;\alpha,\beta)\alpha*_{p}\beta,$
where $\mathrm{sign}(p;\alpha,\beta)$ is the local intersection number of the velocity vectors of $\alpha$ and $\beta$ at $p$ with respect to the orientation of the surface, and $\alpha*_{p}\beta$ is the free loop obtained by first traversing $\alpha$ starting at $p$ then $\beta$.
Definition 2.2.
Assume $n\geq 0$. We define the Kawazumi–Kuno action $\sigma\colon|\mathbb{K}\pi|\to\operatorname{End}_{\mathbb{K}}(\mathbb{K}\pi)$ by, for generically immersed free loops $\alpha$ and $\beta$ on $\Sigma$,
$\sigma(\alpha)(\beta)=\sum_{p\in\alpha\cap\beta}\mathrm{sign}(p;\alpha,\beta)\alpha*_{p}\beta.$
Definition 2.3.
We define the Turaev cobracket $\delta\colon|\mathbb{K}\pi/\mathbb{K}1|\to|\mathbb{K}\pi/\mathbb{K}1|^{\otimes 2}$ by, for a generically immersed free loop $\alpha\colon[0,1]/\{0,1\}\to\Sigma$,
$\displaystyle\delta(\alpha)=\sum_{\begin{subarray}{c}t_{1}\neq t_{2}\in[0,1]\\ \alpha(t_{1})=\alpha(t_{2})\end{subarray}}\mathrm{sign}(\alpha;t_{1},t_{2})\,\alpha|_{[t_{1},t_{2}]}\otimes\alpha|_{[t_{2},t_{1}]}\,,$
where $\mathrm{sign}(\alpha;t_{1},t_{2})$ is the local intersection number of the velocity vectors $\dot{\alpha}(t_{1})$ and $\dot{\alpha}(t_{2})$ with respect to the orientation of $\Sigma$.
If $\Sigma$ admits a framing $\mathsf{fr}$ (i.e., a smooth non-vanishing vector field), we define the framed version $\delta^{\mathsf{fr}}\colon|\mathbb{K}\pi|\to|\mathbb{K}\pi|^{\otimes 2}$ by the same formula but taking a rotation-free representative of $\alpha$.
Theorem 2.4 ([Gol86],[Tur91],[AKKN23]).
The $\mathbb{K}$-linear maps [,[$\cdot$,$\cdot$][ ⋅ , ⋅ ], $\delta$ and $\delta^{\mathsf{fr}}$ are well-defined. The triples $(|\mathbb{K}\pi|,[\cdot,\cdot],\delta^{\mathsf{fr}})$ and $(|\mathbb{K}\pi/\mathbb{K}1|,[\cdot,\cdot],\delta)$ are Lie bialgebras.
Theorem 2.5 ([KK14]).
The map $\sigma$ is well-defined and takes its values in the space of derivations: $\sigma\colon|\mathbb{K}\pi|\to\operatorname{Der}_{\mathbb{K}}(\mathbb{K}\pi)$.
Next, we recall a weight filtration on $\mathbb{K}\pi$. Note that a (decreasing) filtration on a vector space can be pulled back along any linear map and pushed out along a surjection.
Definition 2.6.
Assume $n\geq 0$.
  • Let $\mathcal{C}=(\alpha_{i},\beta_{i},\gamma_{j})_{1\leq i\leq g,1\leq j\leq n}$ be a free-generating system of $\pi$ so that $\alpha_{i}$ and $\beta_{i}$ form a genus pair, $\gamma_{j}$ is a boundary loop representing $\partial_{j}\Sigma$ and
    $\alpha_{1}\beta_{1}\alpha_{1}^{-1}\beta_{1}^{-1}\cdots\alpha_{g}\beta_{g}\alpha_{g}^{-1}\beta_{g}^{-1}\gamma_{1}\cdots\gamma_{n}$
    represents a boundary loop based at $*\in\partial_{0}\Sigma$ (see Figure 2 of [Tan24a]). Denote by $(x_{i},y_{i},z_{j})_{1\leq i\leq g,1\leq j\leq n}$ the corresponding basis of $H=H_{1}(\Sigma;\mathbb{K})$.
  • $\hat{T}(H)\cong\mathbb{K}\langle\!\langle x_{i},y_{i},z_{j}\rangle\!\rangle_{1\leq i\leq g,1\leq j\leq n}$ is the completed free associative algebra over $H$ with respect to the weight grading on $H$ defined by $\mathrm{wt}(x_{i})=\mathrm{wt}(y_{i})=1$ and $\mathrm{wt}(z_{j})=2$.
  • Consider a morphism of $\mathbb{K}$-algebras
    $\theta_{\mathrm{exp}}\colon\mathbb{K}\pi\to\hat{T}(H)\colon\alpha_{i}\mapsto e^{x_{i}},\beta_{i}\mapsto e^{y_{i}}\textrm{ and }\gamma_{j}\mapsto e^{z_{j}}.$
    Define a filtration on $\mathbb{K}\pi$ by the pull-back of the weight filtration by $\theta_{\mathrm{exp}}$. This induces an isomorphism of Hopf algebras on the completion $\widehat{\mathbb{K}\pi}$.
  • Define filtrations on $|\mathbb{K}\pi|$ and $|\mathbb{K}\pi/\mathbb{K}1|$ as the push-out by the natural surjections $\mathbb{K}\pi\twoheadrightarrow|\mathbb{K}\pi|\twoheadrightarrow|\mathbb{K}\pi/\mathbb{K}1|$.
  • A filtration on $\operatorname{Der}_{\mathbb{K}}(\mathbb{K}\pi)$ is the induced one: denoting the filtration degree $\geq k$ part of $\mathbb{K}\pi$ by $F^{k}\mathbb{K}\pi$, $u\colon\mathbb{K}\pi\to\mathbb{K}\pi$ has filtration degree $\geq p$ if and only if $u(F^{k}\mathbb{K}\pi)\subset F^{k+p}\mathbb{K}\pi$ for all $k$.
Definition 2.7.
For $n=-1$, a filtration on $|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ is defined by the push-out along $|\mathbb{K}\pi_{1}(\Sigma_{g,1})|\twoheadrightarrow|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ by capping the unique boundary component by a disk. The rest is analogously defined.
Proposition 2.8.
The $\mathbb{K}$-linear maps [,[$\cdot$,$\cdot$][ ⋅ , ⋅ ], $\sigma$, $\delta$ and $\delta^{\mathsf{fr}}$ are compatible with the filtrations. They all have the filtration degree $(-2)$.
For the proof, see Section 3 of [AKKN23] and references there. Following them, we denote their associated gradeds by $[\cdot,\cdot]_{\mathrm{gr}}\colon|\hat{T}(H)|^{\otimes 2}\to|\hat{T}(H)|$, $\sigma_{\mathrm{gr}}\colon|\hat{T}(H)|\to\operatorname{Der}_{\mathbb{K}}(\hat{T}(H))$, and so on.
3. The Formality and the Kashiwara–Vergne Problem
The formality problem for the Goldman–Turaev Lie algebra asks if it is isomorphic to its associated graded as a Lie bialgebra and, if so, to determine the set of such isomorphisms. In the case of a surface with non-empty boundary, it is completely solved in [AKKN23]: the set SolKV of such isomorphisms is characterised by two equations (KVI) and (KVII). In the special case of $\Sigma=\Sigma_{0,3}$, the problem is, surprisingly, equivalent to the Kashiwara–Vergne problem in Lie theory, and the automorphism group of the Lie bialgebra and that of the associated graded are given by the KV and KRV groups. This section summarises definitions of these spaces and results in [AKKN23].
First, let us recall the definition of some spaces we need. Fix a connected oriented compact surface $\Sigma$ of genus $g$ with $(n+1)$ boundary components with $n\geq 0$ equipped with a framing $\mathsf{fr}$. Take the free-generating system $(\alpha_{i},\beta_{i},\gamma_{j})$ of $\pi$ and the basis $(x_{i},y_{i},z_{j})$ of $H$ as in the last section. We have the following spaces:
  • $\hat{L}(H)=L(\!(x_{i},y_{i},z_{j})\!)_{1\leq i\leq g,1\leq j\leq n}$ is the completed free Lie algebra over $H$, so that $\hat{T}(H)$ is identified with the (completed) universal enveloping algebra $U\!\hat{L}(H)$,
[…]
  • $\operatorname{Aut}^{+}(\hat{L}(H))=\operatorname{exp}(\operatorname{Der}^{+}(\hat{L}(H)))$,
  • the space of tangential derivations:
    $\operatorname{tDer}^{+}(\hat{L}(H))=\{\tilde{u}=(u;u_{1},\dotsc,u_{n}):u\in\operatorname{Der}^{+}(\hat{L}(H)),\,u_{i}\in\hat{L}(H),\,u(z_{j})=[z_{j},u_{j}]\},$
  • the space of tangential automorphisms:
    $\displaystyle\operatorname{tAut}^{+}(\hat{L}(H))$ $\displaystyle=\operatorname{exp}(\operatorname{tDer}^{+}(\hat{L}(H)))$
    =$\displaystyle=\{\tilde{G}=(G;g_{1},\dotsc,g_{n}):G\in\operatorname{Aut}^{+}(\hat{L}(H)),\,g_{j}\in\operatorname{exp}(\hat{L}(H)),\,G(z_{j})=g_{j}^{-1}z_{j}g_{j}\},$
and some elements:
  • $\xi=\log\left(\prod_{i}(e^{x_{i}}e^{y_{i}}e^{-x_{i}}e^{-y_{i}})\prod_{j}e^{z_{j}}\right)\in\hat{L}(H)$,
  • $\omega=\sum_{i}[x_{i},y_{i}]+\sum_{j}z_{j}\in\hat{L}(H)$,
  • $r(s)=\log\left(\frac{e^{s}-1}{s}\right)\in s\mathbb{K}[[s]]$,
  • $\mathbf{r}=\sum_{i}|r(x_{i})+r(y_{i})|,\;\mathbf{r}^{\prime}=\sum_{i}|r(x_{i})+r(y_{i})|+\sum_{j}|r(z_{j})|\in|\hat{T}(H)|$,
  • $a_{i}=\operatorname{rot}^{\mathsf{fr}}(\alpha_{i}),\,b_{i}=\operatorname{rot}^{\mathsf{fr}}(\beta_{i}),\,c_{j}=\operatorname{rot}^{\mathsf{fr}}(\gamma_{j})\in\mathbb{Z}$, and $\mathbf{p}^{\mathsf{fr}}=\sum_{i}|a_{i}y_{i}-b_{i}x_{i}|\in|\hat{T}(H)|$.
Now recall the standard divergence for a free Lie algebra and the integration of a 11 11-cocycle.
Definition 3.1.
  • $\hat{T}(H)$ is regarded as a $U\!\hat{L}(H)$-module by the left multiplication. For $w=x_{i},y_{i},z_{j}$, $d_{w}\colon\hat{L}(H)\to\hat{T}(H)$ is a continuous Lie algebra 11 11-cocycle specified by $d_{w}(w^{\prime})=\delta_{ww^{\prime}}$ for $w^{\prime}=x_{i},y_{i},z_{j}$ using Kronecker’s delta.
  • We define the single divergence $\operatorname{div}_{x,y,z}\colon\operatorname{Der}(\hat{L}(H))\to|\hat{T}(H)|$ by
    $\operatorname{div}_{x,y,z}(u)=\sum_{w=x_{i},y_{i},z_{j}}|d_{w}u(w)|.$
    This is extended to $\operatorname{tDer}(\hat{L}(H))$ by the composition
    $\operatorname{div}_{x,y,z}\colon\operatorname{tDer}(\hat{L}(H))\to\operatorname{Der}(\hat{L}(H))\xrightarrow{\operatorname{div}_{x,y,z}}|\hat{T}(H)|.$
    The single divergence is itself a Lie algebra 11 11-cocycle.
  • For a pro-nilpotent Lie algebra $\mathfrak{g}$, a continuous $\mathfrak{g}$-module $V$ and a 11 11-cocycle $\psi\colon\mathfrak{g}\to V$, its integration is a group 11 11-cocycle $\Psi\colon\exp(\mathfrak{g})\to V$ given by, for $u\in\mathfrak{g}$,
    $\Psi(e^{u})=\frac{e^{u}-1}{u}\cdot\psi(u).$
    The correspondence $\psi\mapsto\Psi$ is $\mathbb{K}$-linear. For the details, see Appendix A of [AKKN23].
  • Since $\operatorname{tDer}^{+}(\hat{L}(H))$ is pro-nilpotent, we denote the integration of $\operatorname{div}_{x,y,z}\colon\operatorname{tDer}^{+}(\hat{L}(H))\to|\hat{T}(H)|$ by
    $\mathsf{j}_{x,y,z}\colon\operatorname{tAut}^{+}(L)\to|\hat{T}(H)|.$
We also have many 11 11-cocycles:
  • $\mathsf{b}^{\mathsf{fr}}\colon\operatorname{tDer}^{+}(\hat{L}(H))\to|\hat{T}(H)|:\tilde{u}\;\mapsto\;\sum_{j}c_{j}|u_{j}|$ and $\mathsf{c}^{\mathsf{fr}}\colon\operatorname{tAut}^{+}(\hat{L}(H))\to|\hat{T}(H)|$ its integration,
  • $\operatorname{div}^{\mathsf{fr}}\colon\operatorname{tDer}^{+}(\hat{L}(H))\to|\hat{T}(H)|:\tilde{u}\;\mapsto\;\operatorname{div}_{x,y,z}(u)-\mathsf{b}^{\mathsf{fr}}(\tilde{u})+u(\mathbf{r}-\mathbf{p}^{\mathsf{fr}})$ and $\mathsf{j}^{\mathsf{fr}}\colon\operatorname{tAut}^{+}(\hat{L}(H))\to|\hat{T}(H)|$ its integration, and
  • $\operatorname{div}^{\mathsf{fr}}_{\mathrm{gr}}\colon\operatorname{tDer}^{+}(\hat{L}(H))\to|\hat{T}(H)|:\tilde{u}\;\mapsto\;\operatorname{div}_{x,y,z}(u)-\mathsf{b}^{\mathsf{fr}}(\tilde{u})$ and $\mathsf{j}^{\mathsf{fr}}_{\mathrm{gr}}\colon\operatorname{tAut}^{+}(\hat{L}(H))\to|\hat{T}(H)|$ its integration.
Finally, we recall the definition of the KV groups and associators.
Definition 3.2.
For $g,n\geq 0$ and a framing $\mathsf{fr}$, the Kashiwara–Vergne group $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}$, the graded version $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$, and the set of the Kashiwara–Vergne associators $\mathrm{SolKV}^{\mathsf{fr}}_{(g,n+1)}$ is defined by the followings:
  • $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}\tilde{G}\in\operatorname{tAut}^{+}(\hat{L}(H)):G(\xi)=\xi,\,\mathsf{j}^{\mathsf{fr}}(\tilde{G})\in\Big{|}\sum_{j}z_{j}\mathbb{K}[[z_{j}]]+\xi^{2}\mathbb{K}[[\xi]]\Big{|}\Big{\}}$,
  • $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}\tilde{G}\in\operatorname{tAut}^{+}(\hat{L}(H)):G(\omega)=\omega,\,\mathsf{j}^{\mathsf{fr}}_{\mathrm{gr}}(\tilde{G})\in\Big{|}\sum_{j}z_{j}\mathbb{K}[[z_{j}]]+\omega^{2}\mathbb{K}[[\omega]]\Big{|}\Big{\}}$, and
  • $\mathrm{SolKV}^{\mathsf{fr}}_{(g,n+1)}=\{\tilde{G}\in\operatorname{tAut}^{+}(\hat{L}(H)):G(\omega)=\xi,\,\mathsf{j}^{\mathsf{fr}}_{\mathrm{gr}}(\tilde{G})-\mathbf{r}+\mathbf{p}^{\mathsf{fr}}\in\Big{|}\sum_{j}z_{j}\mathbb{K}[[z_{j}]]+\xi^{2}\mathbb{K}[[\xi]]\Big{|}\Big{\}}$.
The groups $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}$ and $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$ are pro-unipotent since $\operatorname{tAut}^{+}(\hat{L}(H))$ is. The corresponding Lie algebras are given by the following:
  • $\mathfrak{kv}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}\tilde{g}\in\operatorname{tDer}^{+}(\hat{L}(H)):g(\xi)=0,\,\operatorname{div}^{\mathsf{fr}}(\tilde{g})\in\Big{|}\sum_{j}z_{j}\mathbb{K}[[z_{j}]]+\xi^{2}\mathbb{K}[[\xi]]\Big{|}\Big{\}}$, and
  • $\mathfrak{krv}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}\tilde{g}\in\operatorname{tDer}^{+}(\hat{L}(H)):g(\omega)=0,\,\operatorname{div}^{\mathsf{fr}}_{\mathrm{gr}}(\tilde{g})\in\Big{|}\sum_{j}z_{j}\mathbb{K}[[z_{j}]]+\omega^{2}\mathbb{K}[[\omega]]\Big{|}\Big{\}}$.
One of the main results in [AKKN23] is the following.
Theorem 3.3 ([AKKN23], Theorem 6.27).
For $\Sigma=\Sigma_{g,n+1}$ with $n\geq 0$, an isomorphism of filtered Hopf algebras $\theta\colon\widehat{\mathbb{K}\pi}\to\hat{T}(H)$ with $\mathrm{gr}(\theta)=\operatorname{id}$ gives a solution to the formality problem if and only if $\theta\circ\theta_{\mathrm{exp}}^{-1}$ lifts to an element in $\mathrm{SolKV}_{(g,n+1)}$ up to conjugation by an element of $\operatorname{Aut}^{+}(\hat{L}(H))$.
SolKV is a (two-sided) torsor over KV and KRV, which is apparent from their defining equations.
4. Reformulation of KV Groups
In this section, we reformulate the defining equations of the KV and KRV groups using certain connections in non-commutative geometry. We use notations in [Tan24b] and [Tan24a]. In Section 6, we will deal with the case of closed surfaces based on observations made here.
4.1 Tangential Derivations
We start with tangential derivations, which are best understood in the setting of linear categories. Let $g,n\geq 0$ and denote $(n+1)$ boundary components of $\Sigma_{g,n+1}$ by $\partial_{0}\Sigma,\dotsc,\partial_{n}\Sigma$. Take one base point $*_{j}$ for each $j$ and put $V=\{*_{j}\}_{0\leq j\leq n}$. Now set $\mathscr{G}=\pi_{1}(\Sigma_{g,n+1},V)$, the fundamental groupoid of $\Sigma_{g,n+1}$ with base points $V$.
Definition 4.1.
Let $(\gamma_{j})_{j}$ be as above: $\gamma_{j}$ is a simple loop based at $*_{0}$ representing the $j$-th boundary. We define
$\operatorname{tDer}(\mathbb{K}\pi)=\{(u,u_{1},\dotsc,u_{n})\in\operatorname{Der}_{\mathbb{K}}(\mathbb{K}\pi)\times(\mathbb{K}\pi)^{n}:u(\gamma_{j})=[\gamma_{j},u_{j}]\}.$
Definition 4.2.
Denote the homotopy class of a simple loop based at $*_{j}$ representing $\partial_{j}\Sigma$ abusively by $\partial_{j}\Sigma$. We define $\operatorname{Der}_{\partial}(\mathbb{K}\mathscr{G})$ to be the space of derivations $f$ on a $\mathbb{K}$-linear category $\mathbb{K}\mathscr{G}$ with $f(\partial_{j}\Sigma)=0$ for all $j$.
Proposition 4.3.
We have an identification $\operatorname{tDer}(\mathbb{K}\pi)\cong\operatorname{Der}_{\partial}(\mathbb{K}\mathscr{G})$.
Proof. Suppose we are given a tangential derivation $(u,u_{1},\dotsc,u_{n})$. We construct a $\mathbb{K}$-linear category derivation $f$. For a loop $\alpha$ based at $*_{0}$, we set $f(\alpha)=u(\alpha)$. For a path $\delta$ from $*_{j}$ to $*_{0}$ such that $\delta\gamma_{j}\delta^{-1}$ is a $j$-th boundary loop, we set $\delta^{-1}f(\delta)=u_{j}$. This is well-defined: for another such path $\delta^{\prime}$, the loop $\delta^{\prime-1}\delta$ is based at $*_{0}$ and we have $(\delta^{\prime-1}\delta)\gamma_{j}=\gamma_{j}(\delta^{\prime-1}\delta)$. Since $\mathscr{G}$ is a free groupoid, we have $\delta^{\prime-1}\delta=\gamma_{j}^{-m}$ for some $m\in\mathbb{Z}$. We compute
$\displaystyle\delta^{\prime-1}f(\delta^{\prime})$ $\displaystyle=\gamma_{j}^{-m}\delta^{-1}f(\delta\gamma_{j}^{m})$
$\displaystyle=\gamma_{j}^{-m}\delta^{-1}f(\delta)\gamma_{j}^{m}+\gamma_{j}^{-m}f(\gamma_{j}^{m})$
$\displaystyle=\gamma_{j}^{-m}u_{j}\gamma_{j}^{m}+\sum_{1\leq i\leq m}\gamma_{j}^{-m+i-1}f(\gamma_{j})\gamma_{j}^{m-i}$
$\displaystyle=\gamma_{j}^{-m}u_{j}\gamma_{j}^{m}+\sum_{1\leq i\leq m}\gamma_{j}^{-m+i-1}(\gamma_{j}u_{j}-u_{j}\gamma_{j})\gamma_{j}^{m-i}$
$\displaystyle=\gamma_{j}^{-m}u_{j}\gamma_{j}^{m}-\gamma_{j}^{-m}u_{j}\gamma_{j}^{m}+u_{j}$
$\displaystyle=u_{j}.$
This shows the well-definedness. We also have
$\displaystyle f(\partial_{j}\Sigma)$ $\displaystyle=f(\delta\gamma_{j}\delta^{-1})$
$\displaystyle=f(\delta)\gamma_{j}\delta^{-1}+\delta(\gamma_{j}u_{j}-u_{j}\gamma_{j})\delta^{-1}-\delta\gamma_{j}\delta^{-1}f(\delta)\delta^{-1}$
$\displaystyle=\delta u_{j}\gamma_{j}\delta^{-1}+\delta(\gamma_{j}u_{j}-u_{j}\gamma_{j})\delta^{-1}-\delta\gamma_{j}u_{j}\delta^{-1}$
$\displaystyle=0.$
In the opposite direction, given a derivation $f$ on $\mathbb{K}\mathscr{G}$ with $f(\partial_{j}\Sigma)=0$, it automatically yields a derivation $u$ by restricting $f$ to $\mathbb{K}\pi=\mathbb{K}\pi_{1}(\Sigma,*_{0})\subset\mathbb{K}\mathscr{G}$. We set $u_{j}=\delta^{-1}f(\delta)$. This yields a bijection.∎
Definition 4.4.
Let $\hat{\mathscr{T}}=\hat{T}(H)\langle\delta_{j},\delta_{j}^{-1}\rangle_{1\leq j\leq n}$ be the $\mathbb{K}$-linear category obtained by formally adjoining the morphisms $\delta_{j}$ and $\delta_{j}^{-1}$ from $*_{j}$ to $*_{0}$ to the completed algebra $\hat{T}(H)$ which is the space of endomorphisms at $*_{0}$. A morphism of $\mathbb{K}$-linear categories
$\theta_{\mathrm{exp}}\colon\mathbb{K}\mathscr{G}\to\hat{\mathscr{T}},$
is given as the unique extension with $\theta_{\mathrm{exp}}(\delta_{j})=\delta_{j}$. The element $\delta_{j}$ is defined to have degree 0. A filtration on $\mathbb{K}\mathscr{G}$ is given by the pull-back by $\theta_{\mathrm{exp}}$.
4.2 Connections on Modules
Next, we define a left $\mathbb{K}\mathscr{G}$-module $\mathscr{N}$ and a connection on $\mathscr{N}$, which eventually give the divergence maps in the last section. Recall that, in [Tan24b], the single divergence is based on a left module over a Hopf algebra $A$, from which we obtain a bimodule by a functor $\Phi_{A}$. In the definition, we used $\mathbb{K}$ as a ground field and also as a left $A$-module via the augmentation $\varepsilon\colon A\to\mathbb{K}$. In the setting of linear categories, we have a notion of Hopf groupoids, and these two $\mathbb{K}$’s become different spaces, as we shall see below. We freely use the correspondence on modules between a linear category and its category algebra.
Definition 4.5 ([Fre17], 9.0.2).
A Hopf groupoid over $\mathbb{K}$ is a small category $\mathscr{A}$ enriched over the monoidal category of counital coassociative $\mathbb{K}$-coalgebras together with the anti-homomorphism $\varsigma\colon\mathscr{A}\to\mathscr{A}^{\mathrm{op}}$ of $\mathbb{K}$-linear categories called the antipode, satisfying the Hopf relations.
Definition 4.6.
Let $\mathscr{A}$ be a $\mathbb{K}$-Hopf groupoid. A Hopf derivation of $\mathscr{A}$ is a derivation $f\colon\mathscr{A}\to\mathscr{A}$ in the sense of Definition 4.1 in [Tan24a], which is also a coderivation on each Hom-space. We denote by $\operatorname{Der}_{\mathrm{Hopf}}(\mathscr{A})$ the space of Hopf derivations.
Definition 4.7.
A $\mathbb{K}$-Hopf groupoid structure of the groupoid algebra $\mathbb{K}\mathscr{G}$ is given by the following: a counital coassociative (cocommutative) coalgebra structure on the Hom-spaces given by, for $v,w\in\mathrm{Ob}(\mathscr{G})=V$ and $x\in\mathscr{G}(v,w)$,
$\Delta\colon\mathbb{K}\mathscr{G}(v,w)\to\mathbb{K}\mathscr{G}(v,w)\otimes\mathbb{K}\mathscr{G}(v,w):x\mapsto x\otimes x$
and
$\varepsilon\colon\mathbb{K}\mathscr{G}(v,w)\to\mathbb{K}:x\mapsto 1,$
and the antipode is given by
$\varsigma\colon\mathbb{K}\mathscr{G}(v,w)\to\mathbb{K}\mathscr{G}(w,v):x\to x^{-1}.$
The completion $\widehat{\mathbb{K}\mathscr{G}}$ and its associated graded $\hat{\mathscr{T}}$ are automatically (complete) Hopf groupoids. We set
$\displaystyle\operatorname{Der}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}})$ =$\displaystyle=\{f\in\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}}):f(\partial_{j}\Sigma)=0\}\textrm{ and }$
$\displaystyle\operatorname{Der}^{+}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}})$ =$\displaystyle=\{f\in\operatorname{Der}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}}):\mathrm{gr}(f)=\operatorname{id}\}.$
Remark 4.8.
The completed free Lie algebra $\hat{L}(H)$ in the last section is the primitive part of $\hat{T}(H)$, and $\operatorname{Der}(\hat{L}(H))$ is canonically isomorphic to the space of (continuous) Hopf derivations $\operatorname{Der}_{\mathrm{Hopf}}(\hat{T}(H))$. Since the notion of the primitive part behaves poorly in the setting of multi-objects, we stick to the perspective of Hopf algebras: we use $\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}})$ as a multi-object analogue of $\operatorname{Der}(\hat{L}(H))$.
Definition 4.9.
Let $\mathscr{A}$ be a $\mathbb{K}$-Hopf groupoid.
  • We define the $\mathbb{K}$-linear category $\mathscr{E}$ by $\mathrm{Ob}(\mathscr{E})=\mathrm{Ob}(\mathscr{A})$ and $\mathscr{E}(v,w)=\mathbb{K}$ for $v,w\in\mathrm{Ob}(\mathscr{A})$ with the composition map
    $\mathscr{E}(u,v)\otimes\mathscr{E}(v,w)=\mathbb{K}\otimes\mathbb{K}\xrightarrow{\cong}\mathbb{K}=\mathscr{E}(u,w).$
    Denote the unit $1\in\mathbb{K}=\mathscr{E}(v,w)$ by $e_{vw}$. Then $\varepsilon$ above assembles into a morphism of $\mathbb{K}$-linear categories $\varepsilon\colon\mathscr{A}\to\mathscr{E}$ which reads $\varepsilon(x)=e_{vw}$ in this notation.
  • Let $\mathbb{K}\mathrm{Ob}(\mathscr{A})$ be a left $\mathscr{E}$-module with $\mathbb{K}\mathrm{Ob}(\mathscr{A})(v)=\mathbb{K}v$, the one-dimensional space generated by the symbol $v$. The action of $\mathscr{E}$ is given explicitly by $e_{vw}\cdot w=v$.
  • We put $s\colon\mathscr{A}\to\mathbb{K}\mathrm{Ob}(\mathscr{A})\colon x\mapsto\sum_{v\in\mathrm{Ob}(\mathscr{A})}\varepsilon(x)\cdot v$. We call it the source map since for $x\in\mathscr{A}(v,w)$ with $\varepsilon(x)=e_{vw}$, we have $s(x)=v$. This is a left module map over $\varepsilon$.
  • Recall that $V=\mathrm{Ob}(\mathbb{K}\mathscr{G})$. Regarding $\mathbb{K}V$ as a $\mathbb{K}\mathscr{G}$-module via $\varepsilon$, we define another $\mathbb{K}\mathscr{G}$-module by $\mathscr{N}=\operatorname{Ker}(s\colon\mathbb{K}\mathscr{G}\to\mathbb{K}V)$.
Remark 4.10.
[…]
Lemma 4.11.
$\mathscr{N}$ is dualisable over $\mathbb{K}\mathscr{G}$.
Proof. We have an identification as $\mathbb{K}\mathscr{G}$-modules:
$\mathscr{N}\cong\bigoplus_{c=\alpha_{i},\beta_{i},\gamma_{j}}\mathbb{K}\mathscr{G}(\,\cdot\,,*_{0})(1_{0}-c)\oplus\bigoplus_{1\leq j\leq n}\mathbb{K}\mathscr{G}(\,\cdot\,,*_{j})(1_{j}-\delta_{j}).$
The proof is analogous to Proposition 3.8 in [Tan24b] and Lemma 5.6 in [Tan24a].∎
Now we define a connection on $\mathscr{N}$. For the definition of a connection on a module over a linear category, see Definition 4.8 of [Tan24a].
Definition 4.12.
Take a free-generating system of $\mathscr{G}$ as in Figure 2 of [Tan24a]. We define a connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}\colon\mathscr{N}\to\Omega^{1}\mathbb{K}\mathscr{G}\otimes_{\mathbb{K}\mathscr{G}}\mathscr{N}$ by
$\displaystyle\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}(1_{0}-\alpha_{i})=a_{i}(d\beta_{i})\beta_{i}^{-1}\otimes(1_{0}-\alpha_{i}),$
$\displaystyle\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}(1_{0}-\beta_{i})=-b_{i}(d\alpha_{i})\alpha_{i}^{-1}\otimes(1_{0}-\beta_{i}),$
$\displaystyle\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}(1_{0}-\gamma_{j})=0,\textrm{ and}$
$\displaystyle\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}(1_{j}-\delta_{j})=c_{j}(d\delta_{j})\delta_{j}^{-1}\otimes(1_{j}-\delta_{j}).$
Extend $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ to the completion $\hat{\mathscr{N}}$ by continuity, which is a $\widehat{\mathbb{K}\mathscr{G}}$-module. A derivation action of $f\in\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}})$ on the module $\hat{\mathscr{N}}$ is given by a restriction $f|_{\operatorname{Ker}(s)}$.
Definition 4.13.
Let $\hat{\mathscr{N}}_{\mathrm{gr}}=\operatorname{Ker}(s\colon\hat{\mathscr{T}}\to\mathbb{K}V)$, which is also dualisable over $\hat{\mathscr{T}}$. We define a connection $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on $\hat{\mathscr{N}}_{\mathrm{gr}}$ by
$\nabla^{\prime}\!\!_{H,\mathsf{fr}}(x_{i})=\nabla^{\prime}\!\!_{H,\mathsf{fr}}(y_{i})=\nabla^{\prime}\!\!_{H,\mathsf{fr}}(z_{j})=0\;\textrm{ and }\;\nabla^{\prime}\!\!_{H,\mathsf{fr}}(1_{j}-\delta_{j})=c_{j}(d\delta_{j})\delta_{j}^{-1}\otimes(1_{j}-\delta_{j}).$
This is an extension of the flat connection $\nabla^{\prime}\!\!_{z}$ defined in Section 6 of [Tan24b]. Also this is the associated graded of $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ above.
Lemma 4.14.
We have $\operatorname{div}^{\mathsf{fr}}=-\operatorname{Div}^{\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}}$ as maps $\operatorname{Der}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}})\to|\widehat{\mathbb{K}\mathscr{G}}|$, and $\operatorname{div}^{\mathsf{fr}}_{\mathrm{gr}}=-\operatorname{Div}^{\nabla^{\prime}\!\!_{H,\mathsf{fr}}}$ as maps $\operatorname{Der}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}})\to|\hat{\mathscr{T}}|$.
Before the proof, we look at a counterpart of the connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ on a $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$. For a $\mathbb{K}$-algebra $A$, we denote the copy of an element $a\in A$ in $A^{\mathrm{op}}$ by $\bar{a}$. We denote by $A^{\mathrm{e}}$ the enveloping algebra of $A$.
Definition 4.15.
We define a connection $\nabla\!_{\mathcal{C},\mathsf{fr}}\colon\Omega^{1}\mathbb{K}\mathscr{G}\to\Omega^{1}\mathbb{K}\mathscr{G}^{\mathrm{e}}\otimes_{\mathbb{K}\mathscr{G}^{\mathrm{e}}}\Omega^{1}\mathbb{K}\mathscr{G}$ on a $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$ by
$\displaystyle\nabla\!_{\mathcal{C},\mathsf{fr}}\!\left((d\alpha_{i})\alpha_{i}^{-1}\right)=a_{i}[(d\beta_{i})\beta_{i}^{-1}-\overline{(d\beta_{i})\beta_{i}^{-1}}]\otimes(d\alpha_{i})\alpha_{i}^{-1},$
$\displaystyle\nabla\!_{\mathcal{C},\mathsf{fr}}\!\left((d\beta_{i})\beta_{i}^{-1}\right)=-b_{i}[(d\alpha_{i})\alpha_{i}^{-1}-\overline{(d\alpha_{i})\alpha_{i}^{-1}}]\otimes(d\beta_{i})\beta_{i}^{-1}$
$\displaystyle\nabla\!_{\mathcal{C},\mathsf{fr}}\!\left((d\gamma_{j})\gamma_{j}^{-1}\right)=0,\textrm{ and}$
$\displaystyle\nabla\!_{\mathcal{C},\mathsf{fr}}\!\left((d\delta_{j})\delta_{j}^{-1}\right)=c_{j}[(d\delta_{j})\delta_{j}^{-1}-\overline{(d\delta_{j})\delta_{j}^{-1}}]\otimes(d\delta_{j})\delta_{j}^{-1}.$
Theorem 4.16.
Let $\sigma\colon|\mathbb{K}\mathscr{G}|\to\operatorname{Der}_{\partial}(\mathbb{K}\mathscr{G})$ be the groupoid version of the Kawazumi–Kuno action. For any framing $\mathsf{fr}$, the framed version of the Turaev cobracket $\delta^{\mathsf{fr}}$ is equal to the composition $-\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}\circ\,\sigma$.
Proof. Since the natural map $|\mathbb{K}\mathscr{G}|\to|\widehat{\mathbb{K}\mathscr{G}}|$ is injective, all the computations can be done in the completion, which is further identified with $|\hat{\mathscr{T}}|$ by $\theta_{\mathrm{exp}}$. For $f\in\operatorname{Der}_{\partial}(\mathbb{K}\mathscr{G})$, the associated divergence is given by
$\displaystyle-\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}(f)$ $\displaystyle=-\operatorname{Tr}(i_{f}\nabla\!_{\mathcal{C},\mathsf{fr}}-L_{f})$
$\displaystyle=\sum_{1\leq i\leq g}-|a_{i}(f(\beta_{i})\beta_{i}^{-1}-\overline{f(\beta_{i})\beta_{i}^{-1}})|+|b_{i}(f(\alpha_{i})\alpha_{i}^{-1}-\overline{f(\alpha_{i})\alpha_{i}^{-1}})|$
$\displaystyle\qquad-\sum_{1\leq j\leq n}|c_{j}(f(\delta_{j})\delta_{j}^{-1}-\overline{f(\delta_{j})\delta_{j}^{-1}})|+\sum_{c\in\mathcal{C}}|\partial_{c}(f(c))-1\otimes c^{-1}f(c)|$
$\displaystyle=\sum_{1\leq i\leq g}a_{i}|1\wedge f(\beta_{i})\beta_{i}^{-1}|-b_{i}|1\wedge f(\alpha_{i})\alpha_{i}^{-1}|$
$\displaystyle\qquad+\sum_{1\leq j\leq n}c_{j}|1\wedge f(\delta_{j})\delta_{j}^{-1}|+\sum_{c\in\mathcal{C}}|\partial_{c}(f(c))-1\otimes c^{-1}f(c)|$
$\displaystyle=1\wedge f(\mathbf{p}^{\mathsf{fr}})+1\wedge\mathsf{b}^{\mathsf{fr}}(f)+\sum_{c=\alpha_{i},\beta_{i},\gamma_{j}}|\partial_{c}(f(c))-1\otimes c^{-1}f(c)|+\sum_{1\leq j\leq n}|\partial_{\delta_{j}}(f(\delta_{j}))-1\otimes\delta_{j}^{-1}f(\delta_{j})|.$
The last term is equal to zero since we have
$\partial_{\delta_{j}}(f(\delta_{j}))-1\otimes\delta_{j}^{-1}f(\delta_{j})=\partial_{\delta_{j}}(\delta_{j}u_{j})-1\otimes u_{j}=1\otimes u_{j}-1\otimes u_{j}=0.$
Therefore we have $-\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}=\operatorname{Div}^{\mathsf{fr}}$ on $\operatorname{Der}_{\partial}(\mathbb{K}\mathscr{G})$, where $\operatorname{Div}^{\mathsf{fr}}$ is defined in Definition 5.8 of [AKKN23]. By Theorem 5.16 of [AKKN23], we have $\delta^{\mathsf{fr}}=\operatorname{Div}^{\mathsf{fr}}\circ\,\sigma$. This completes the proof.∎
The above wording “counterpart” is justified by the following.
Lemma 4.17.
The connection $\nabla\!_{\mathcal{C},\mathsf{fr}}$ is induced by $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ in the sense of Definition-Lemma 6.2 in [Tan24b].
Proof. Let $A$ be the category algebra of $\mathbb{K}\mathscr{G}$, and $S\cong\prod_{V}\mathbb{K}$ the corresponding subalgebra generated by the identity morphisms in $\mathbb{K}\mathscr{G}$. The induced connection is only defined in the absolute case $S=\mathbb{K}$ in [Tan24b], so we describe the construction for the general case here. The reader may want to refer to Section 6 there beforehand.
First, the source map corresponds to a left $A$-module map $s\colon A\to\mathbb{K}V$, where the corresponding $A$-module to $\mathbb{K}V$ is denoted again by $\mathbb{K}V$. In addition, $\mathscr{N}$ corresponds to $N=\operatorname{Ker}(s\colon A\to\mathbb{K}V)$. Next, consider the functor
$\Phi_{(A,S)}\colon A\textrm{-}\textbf{Mod}\to A^{\mathrm{e}}\textrm{-}\textbf{Mod}:M\mapsto\Phi_{(A,S)}(M),\;\;\psi\mapsto\psi\otimes\operatorname{id}_{A}\,.$
Here we set $\Phi_{(A,S)}(M)=M\otimes_{S}A$ as $\mathbb{K}$-vector spaces with $M$ regarded as a right $S$-module since $S$ is commutative, and the bimodule structure is given by, for $a,x,y\in A$ and $m\in M$,
$x\cdot(m\otimes a)\cdot y=x^{(1)}m\otimes x^{(2)}ay,$
using the coproduct $\Delta(x)=x^{(1)}\otimes x^{(2)}$. This action is well-defined: for two representatives $1_{v}m\otimes a$ and $m\otimes 1_{v}a$ of the same element in $\Phi_{(A,S)}(M)$, the action of $x\in\mathscr{G}(-,w)$ is computed as, if $w=v$,
$x\cdot(1_{v}m\otimes a)=xm\otimes xa=x\cdot(m\otimes 1_{v}a),$
and if $w\neq v$,
$x\cdot(1_{v}m\otimes a)=0=x\cdot(m\otimes 1_{v}a).$
Then, we have isomorphisms of $A$-bimodules: for $x,y\in\mathscr{G}$ and $v\in V$,
$\displaystyle\Phi_{(A,S)}(A)\cong A\otimes_{S}A$ $\displaystyle:x\otimes y\mapsto x\otimes x^{-1}y\;\textrm{ and}$ (1)
$\displaystyle\Phi_{(A,S)}(\mathbb{K}V)\cong A\hskip 25.5pt$ $\displaystyle:v\otimes x\mapsto 1_{v}x\,.$
The functor $\Phi_{(A,S)}$ is exact since $S$ being separable implies every (left) module over $S$ is projective. By applying it to the short exact sequence
[…]
combined with the two isomorphisms above, we get
$0\to\Phi_{(A,S)}(N)\to A\otimes_{S}A\xrightarrow{\mathrm{mult}}A\to 0,$
so we have $\Phi_{(A,S)}(N)\cong\operatorname{Ker}(A\otimes_{S}A\xrightarrow{\mathrm{mult}}A)=\Omega^{1}_{S}A$. Furthermore, we have a module map
$j_{N}\colon N\to\Phi_{(A,S)}(N):n\mapsto n\otimes 1$
over the twisted coproduct $\tilde{\Delta}=(\operatorname{id}\otimes\,\varsigma)\circ\Delta$ by Definition-Lemma 6.1 of [Tan24b]. The composition of $j_{N}$ with the isomorphism (1) sends $(1_{0}-c)$ to $(dc)c^{-1}$ for $c\in\mathcal{C}$, and $(1_{j}-\delta_{j})$ to $(d\delta_{j})\delta_{j}^{-1}$, respectively. Since the induced connection $\Phi_{(A,S)}(\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})$ is the natural extension of $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ along $j_{N}$, we have $\Phi_{(A,S)}(\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})=\nabla\!_{\mathcal{C},\mathsf{fr}}$ under the identification (1).∎
Proof of Lemma 4.14. Since $\Phi_{(A,S)}(\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})=\nabla\!_{\mathcal{C},\mathsf{fr}}$, we can apply the Hopf groupoid version of Theorem 6.4 in [Tan24b] to conclude $\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}=|\tilde{\Delta}|\circ\operatorname{Div}^{\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}}$ on $\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}})$, where $|\tilde{\Delta}|\colon|\widehat{\mathbb{K}\mathscr{G}}|\to|\widehat{\mathbb{K}\mathscr{G}}|\otimes|\widehat{\mathbb{K}\mathscr{G}^{\mathrm{op}}}|$ is the map induced by the twisted coproduct. Using the augmentation map $|\varepsilon\circ\varsigma|\colon|\widehat{\mathbb{K}\mathscr{G}^{\mathrm{op}}}|\to|\mathscr{E}|\cong\mathbb{K}$, we obtain
$(\operatorname{id}\otimes|\varepsilon\circ\varsigma|)\circ\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}=\operatorname{Div}^{\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}}.$
We also have $-\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}=\operatorname{Div}^{\mathsf{fr}}$ on $\operatorname{Der}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}})$ by the proof of Theorem 4.16, which is further equal to $|\tilde{\Delta}|\circ\operatorname{div}^{\mathsf{fr}}$ by Proposition 4.42 of [AKKN23]. Hence, we have
$-\operatorname{Div}^{\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}}=-(\operatorname{id}\otimes|\varepsilon\circ\varsigma|)\circ\operatorname{Div}^{\nabla\!_{\mathcal{C},\mathsf{fr}}}=(\operatorname{id}\otimes|\varepsilon\circ\varsigma|)\circ|\tilde{\Delta}|\circ\operatorname{div}^{\mathsf{fr}}=\operatorname{div}^{\mathsf{fr}}.$
For $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$, it can be similarly done.∎
4.3 Integration and Connections
Our goal, for the time being, is to explicitly integrate divergences in terms of connections (and this is essentially done in Proposition 4.39 of [AKKN23]). First, we define the adjoint action.
Definition 4.18.
Let $B$ be a $\mathbb{K}$-algebra and $R$ a subalgebra of $B$.
  • An automorphic action on a $B$-module $M$ relative to $R$ is a triple $(\Gamma,\varphi,\rho)$ where $\Gamma$ is a group, and
    $\varphi\colon\Gamma\to\operatorname{Aut}_{R}(B)\;\textrm{ and }\;\rho\colon\Gamma\to\operatorname{Aut}_{\mathbb{K}}(M)$
    are group homomorphisms satisfying
    $\rho(\gamma)(bm)=\varphi(\gamma)(b)\cdot\rho(\gamma)(m).$
    for $\gamma\in\Gamma$, $b\in B$ and $m\in M$. Here $\operatorname{Aut}_{R}(B)$ is the set of algebra automorphisms that restricts to the identity on $R$.
  • Let $\nabla\colon M\to\Omega^{1}_{R}B\otimes_{B}M$ be an $R$-linear connection on $M$ with an automorphic action. We define the adjoint action of $\gamma\in\Gamma$ to $\nabla$ by the composite
    $\operatorname{Ad}_{\gamma}\!\nabla\colon M\xrightarrow{\rho(\gamma^{-1})}M\xrightarrow{\nabla}\Omega^{1}_{R}B\otimes_{B}M\xrightarrow{\varphi(\gamma)\otimes\rho(\gamma)}\Omega^{1}_{R}B\otimes_{B}M\,.$
Recall that the space of de Rham forms $\operatorname{DR}_{R}^{\bullet}\!B$ is the graded cyclic quotient of $\Omega^{\bullet}_{R}B$.
Proposition 4.19.
Let $B$ and $M$ be equipped with a topology, $\mathfrak{d}$ a pro-unipotent Lie algebra and $(\mathfrak{d},\varphi,\rho)$ a continuous derivation action $M$. Let $\nabla$ be a trace-flat $R$-linear continuous connection on $M$, namely, $\operatorname{Tr}(\nabla^{2})=0$ in $\operatorname{DR}_{R}^{2}\!B$. We denote the integration of $\operatorname{Div}^{\nabla}\colon\mathfrak{d}\to|B|$ by $\mathsf{J}^{\nabla}\colon\operatorname{exp}(\mathfrak{d})\to|B|$. Then, for $G\in\operatorname{exp}(\mathfrak{d})$, we have $d(\mathsf{J}^{\nabla}(G))=\operatorname{Tr}((\operatorname{Ad}_{G}-\operatorname{id})\nabla)$ in $\operatorname{DR}_{R}^{1}\!B$.
Proof. First, the derivation action $(\mathfrak{d},\varphi,\rho)$ induces an automorphic action $(\operatorname{exp}(\mathfrak{d}),\varphi,\rho)$, and the map $G\mapsto(\operatorname{Ad}_{G}-\operatorname{id})\nabla$ on the right-hand side is a group 11 11-cocycle. Since the trace map is $\operatorname{Aut}_{R}(B)$-equivariant, the map $G\mapsto\operatorname{Tr}((\operatorname{Ad}_{G}-\operatorname{id})\nabla)$ is also a group 11 11-cocycle. On the other hand, the trace-flatness implies $\operatorname{Div}^{\nabla}$ is a Lie algebra 11 11-cocycle:
$\operatorname{Div}^{\nabla}([f,g])-f(\operatorname{Div}^{\nabla}(g))+g(\operatorname{Div}^{\nabla}(f))=i_{\varphi(f)}i_{\varphi(g)}\operatorname{Tr}(\nabla^{2})$
for $f,g\in\mathfrak{d}$ where $i$ denotes the contraction. The proof is essentially done in Lemma 4.11 of [Tan24b]. Therefore $d(\mathsf{J}^{\nabla})$ is a group 11 11-cocycle. Now, it suffices to show that differentiated Lie algebra 11 11-cocycle satisfies the equation
$d(\operatorname{Div}^{\nabla}(f))=\operatorname{Tr}\left(\left.\frac{d}{d\varepsilon}\right|_{\varepsilon=0}(\operatorname{Ad}_{e^{\varepsilon f}}-\operatorname{id})\nabla\right)$
for $f\in\mathfrak{d}$. It is readily seen that $\left.\frac{d}{d\varepsilon}\right|_{\varepsilon=0}(\operatorname{Ad}_{e^{\varepsilon f}}-\operatorname{id})\nabla$ is equal to the infinitesimal adjoint action
$\operatorname{ad}_{f}\!\nabla:=(L_{f}\otimes\operatorname{id}+\operatorname{id}\otimes\rho(f))\circ\nabla-\nabla\circ\rho(f),$
and that $\operatorname{Tr}(\operatorname{ad}_{f}\!\nabla)=d(\operatorname{Div}^{\nabla}(f))$ by Corollary A.6 of [Tan25]. ∎
With these preparations, we can finally write down the KV groups in terms of connections. We introduce some more notations.
Definition 4.20.
  • Let $\partial\widehat{\mathbb{K}\mathscr{G}}$ be a topological linear subcategory of $\widehat{\mathbb{K}\mathscr{G}}$ generated by boundary loops $\partial_{j}\Sigma$, and $\partial\hat{\mathscr{T}}$ its associated graded. We have identifications $|\partial\widehat{\mathbb{K}\mathscr{G}}|=\Big{|}\sum_{j}\mathbb{K}[[z_{j}]]+\mathbb{K}[[\xi]]\Big{|}$ and $|\partial\hat{\mathscr{T}}|=\Big{|}\sum_{j}\mathbb{K}[[z_{j}]]+\mathbb{K}[[\omega]]\Big{|}$.
  • We set $\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}})=\{G\in\operatorname{Aut}^{+}_{\mathrm{Hopf}}(\hat{\mathscr{T}}):G|_{\partial\hat{\mathscr{T}}}=\operatorname{id}\}$ and
    $\displaystyle\mathrm{Isom}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}},\widehat{\mathbb{K}\mathscr{G}})$ =$\displaystyle=\{G\colon\hat{\mathscr{T}}\to\widehat{\mathbb{K}\mathscr{G}}:\textrm{a complete Hopf groupoid isomorphism, }$,
    $\displaystyle\qquad\quad\mathrm{gr}(G)=\operatorname{id},\textrm{and restricts to }G\colon\partial\hat{\mathscr{T}}\xrightarrow{\cong}\partial\widehat{\mathbb{K}\mathscr{G}}\}.$
Theorem 4.21.
We get the following expression:
  • $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})\in d|\partial\widehat{\mathbb{K}\mathscr{G}}|\Big{\}}$,
  • $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!\!_{H,\mathsf{fr}})\in d|\partial\hat{\mathscr{T}}|\Big{\}}$, and
  • $\mathrm{SolKV}^{\mathsf{fr}}_{(g,n+1)}=\Big{\{}G\in\mathrm{Isom}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}},\widehat{\mathbb{K}\mathscr{G}}):\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})\in d|\partial\widehat{\mathbb{K}\mathscr{G}}|\Big{\}}$.
Proof. Since $\log G$ has degree $\geq 1$, traces in the above expressions have no constant terms. Therefore, combining with $\operatorname{Ker}(d\colon|\hat{\mathscr{T}}|\to\operatorname{DR}^{1}\!\hat{\mathscr{T}})=\mathbb{K}$, we can work in the space of de Rham 11 11-forms without any loss of information. The isomorphisms
$\displaystyle\operatorname{tAut}^{+}(\hat{L}(H))\cap\{G(\xi)=\xi\}$ $\displaystyle\cong\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\widehat{\mathbb{K}\mathscr{G}}),$
$\displaystyle\operatorname{tAut}^{+}(\hat{L}(H))\cap\{G(\omega)=\omega\}$ $\displaystyle\cong\operatorname{Aut}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}}),\textrm{ and}$
$\displaystyle\operatorname{tAut}^{+}(\hat{L}(H))\cap\{G(\omega)=\xi\}$ $\displaystyle\cong\mathrm{Isom}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}},\widehat{\mathbb{K}\mathscr{G}})$
are obtained from Proposition 4.3 together with the isomorphism $\theta_{\mathrm{exp}}$. Next, we can check that two connections $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ and $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ are trace-flat. By Lemma 4.14 and Proposition 4.19, group 11 11-cocycles $d(\mathsf{j}^{\mathsf{fr}}(G))$ and $d(\mathsf{j}^{\mathsf{fr}}_{\mathrm{gr}}(G))$ are equal to $-\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})$ and $-\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!\!_{H,\mathsf{fr}})$, respectively. This completes the proof for $\mathrm{KV}^{\mathsf{fr}}_{(g,n+1)}$ and $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$.
For $\mathrm{SolKV}^{\mathsf{fr}}_{(g,n+1)}$, we first compute
$\displaystyle(\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\operatorname{Ad}_{\theta_{\mathrm{exp}}}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})(1_{0}-e^{x_{i}})$ $\displaystyle=\nabla^{\prime}\!\!_{H,\mathsf{fr}}(1_{0}-e^{x_{i}})-\theta_{\mathrm{exp}}(\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})(1_{0}-\alpha_{i})$
$\displaystyle=d\left(\frac{1_{0}-e^{x_{i}}}{x_{i}}\right)\otimes x_{i}-a_{i}(de^{y_{i}})e^{-y_{i}}\otimes(1_{0}-e^{x_{i}})$
$\displaystyle=\left(d\left(\frac{1_{0}-e^{x_{i}}}{x_{i}}\right)\cdot\frac{x_{i}}{1_{0}-e^{x_{i}}}-a_{i}(de^{y_{i}})e^{-y_{i}}\right)\otimes(1_{0}-e^{x_{i}}),$
whose coefficient is equal to $|dr(x_{i})-a_{i}dy_{i}|$ in $\operatorname{DR}^{1}\hat{\mathscr{T}}$. It is similar for $y_{i}$ and $z_{j}$. For $\delta_{j}$, we have
$(\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\operatorname{Ad}_{\theta_{\mathrm{exp}}}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})(1_{j}-\delta_{j})=0$
by definition. Therefore, we have
$\displaystyle\operatorname{Tr}(\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\operatorname{Ad}_{\theta_{\mathrm{exp}}}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})=\sum_{i}d|r(x_{i})-a_{i}y_{i}+r(y_{i})+b_{i}x_{i}|+\sum_{j}d|r(z_{j})|=d|\mathbf{r}^{\prime}-\mathbf{p}^{\mathsf{fr}}|.$
Now for $G\in\mathrm{Isom}^{+}_{\mathrm{Hopf},\partial}(\hat{\mathscr{T}},\widehat{\mathbb{K}\mathscr{G}})$, we put $F=\theta_{\mathrm{exp}}\circ G\in\operatorname{Aut}^{+}_{\mathrm{Hopf}}(\hat{\mathscr{T}})$. We compute
$\displaystyle\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})$ $\displaystyle=\operatorname{Tr}(\operatorname{Ad}_{\theta_{\mathrm{exp}}^{-1}\circ F}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\operatorname{Ad}_{\theta_{\mathrm{exp}}^{-1}}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}})+\operatorname{Tr}(\operatorname{Ad}_{\theta_{\mathrm{exp}}^{-1}}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})$
$\displaystyle=\theta_{\mathrm{exp}}^{-1}\left[\operatorname{Tr}(\operatorname{Ad}_{F}\!\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\nabla^{\prime}\!\!_{H,\mathsf{fr}})+\operatorname{Tr}(\nabla^{\prime}\!\!_{H,\mathsf{fr}}-\operatorname{Ad}_{\theta_{\mathrm{exp}}}\!\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}})\right]$
$\displaystyle=\theta_{\mathrm{exp}}^{-1}\left[d(\mathsf{J}^{\nabla^{\prime}\!\!_{H,\mathsf{fr}}}(F))+d|\mathbf{r}^{\prime}-\mathbf{p}^{\mathsf{fr}}|\right]$
$\displaystyle=\theta_{\mathrm{exp}}^{-1}\left[-d(\mathsf{j}^{\mathsf{fr}}_{\mathrm{gr}}(F))+d|\mathbf{r}^{\prime}-\mathbf{p}^{\mathsf{fr}}|\right].$
Using $\mathbf{r}-\mathbf{r}^{\prime}\in\partial\widehat{\mathbb{K}\mathscr{G}}$, this is exactly the defining equation of the KV associators. ∎
5. Rewriting Rules
Let $T(H)_{\omega}=T(H)/\langle\omega\rangle$, which is the quotient of (non-completed) $T(H)$ by the two-sided ideal $\langle\omega\rangle$. The purpose of this section is to construct a basis of $|T(H)_{\omega}|$ for the later sections. See Theorem 5.22 for the final result.
A basis is constructed by means of rewriting rules. First of all, consider the following (one-way) rewriting in $|T(H)|$: for any monomial $b$ in $T(H)$,
[…]
Here we put $\omega^{\prime}=\sum_{1\leq i<g}[x_{i},y_{i}]$. This rewriting may not terminate, so we introduce the irregularity to measure the number of steps for the termination.
Definition 5.1.
Let $C$ be the set of all (non-commutative) monomials over $H$ so that $C$ is a basis of $|T(H)|$. For $w\in C$, its irregularity $\mathrm{irr}(w)$ is defined by the following steps. For each appearance $\eta$ of $y_{g}$ in $w$, compute $t_{\eta}\in\mathbb{Z}_{\geq 0}\cup\{+\infty\}$ by the following procedure:
  1. (1)
    Start from the letter $\eta$ in $w$. Set $t_{\eta}=0$;
  2. (2)
    See the next (cyclically adjacent to the right) letter $h$. If $h=x_{g}$, add 11 11 to $t_{\eta}$. If $h=y_{g}$, add 00 0 to $t_{\eta}$. Otherwise, terminate the procedure;
  3. (3)
    Go back to (2).
If this procedure terminates, we define $t_{\eta}$ as the final value we obtained. If not, we set $t_{\eta}$ as the supremum of the values attained during the procedure, either 00 0 or +$\infty$+ ∞. We define the irregularity of $w$ by
$\operatorname{irr}(w)=\sum_{\eta}t_{\eta}\in\mathbb{Z}_{\geq 0}\cup\{+\infty\}.$
If there is no $y_{g}$’s in $w$, the sum is empty and $\operatorname{irr}(w)=0$. For a polynomial $a=\sum_{w\in C}c_{w}w\in|T(H)|$ with $c_{w}\in\mathbb{K}$, we set $\operatorname{irr}(a)=\sup_{w:c_{w}\neq 0}(\operatorname{irr}(w))\in\mathbb{Z}_{\geq 0}\cup\{+\infty\}$.
Example 5.2.
Assume $g\geq 2$.
  • If $w$ is the empty word (corresponding to $|1|\in|T(H)|$), we have $\operatorname{irr}(w)=0$.
  • Consider the word $w=|y_{g}x_{g}x_{1}|$. There is only one $y_{g}$ in $w$, so we set $\eta$ as that. Following the procedure above, we get $t_{\eta}=1$. We have $\operatorname{irr}(w)=1$.
  • Consider the word $w=|y_{g}x_{g}x_{g}y_{g}x_{g}y_{1}|$. There are two $y_{g}$’s in $w$, and each yields $t_{\eta}=3$ and 11 11, respectively. We have $\operatorname{irr}(w)=3+1=4$.
  • Consider the word $w=|y_{g}x_{g}x_{g}|$. There is only one $y_{g}$ in $w$. The above procedure does not terminate, encountering infinitely many $x_{g}$’s throughout. We have $\operatorname{irr}(w)=+\infty$.
  • Consider the word $w=|y_{g}|$. There is only one $y_{g}$ in $w$. The above procedure does not terminate while encountering no $x_{g}$’s. We have $\operatorname{irr}(w)=0$.
Lemma 5.3.
For $w\in C$, we have $\operatorname{irr}(w)=+\infty$ if and only if all of the followings hold:
  1. (1)
    $w$ consists only of $x_{g}$’s and $y_{g}$’s;
  2. (2)
    $w$ contains both $x_{g}$ and $y_{g}$.
Proof. Assume (1) and (2). By condition (1), the procedure above does not terminate for any choice of $\eta$. From (2), we can see that $t_{\eta}=+\infty$.
In the other direction, let us assume the negation of (1). Then, the above procedure terminates, and $t_{\eta}$ is always finite for any $\eta$. If we assume (1) and the negation of (2), the only possible cyclic word is either $|x_{g}^{m}|$ or $|y_{g}^{m}|$ for some $m\geq 0$. In any case, we have $\operatorname{irr}(w)=0<+\infty$.∎
Lemma 5.4.
Rewriting by $\rho$ reduces the irregularity: for $a\in|T(H)|$ with $0<\operatorname{irr}(a)<+\infty$, we can obtain $a^{\prime}\in|T(H)|$ by applying $\rho$ such that $\operatorname{irr}(a^{\prime})<\operatorname{irr}(a)$.
Proof. It suffices to see the case $a=w\in C$. By the assumption $\operatorname{irr}(w)>0$, there is a sequence $y_{g}x_{g}$ in $w$. We put $w=|by_{g}x_{g}|$ for some monomial $b\in T(H)$. By the assumption $\operatorname{irr}(w)<+\infty$, $b$ contains $x_{i}$ or $y_{i}$ for $1\leq i<g$. By applying $\rho$, we obtain a polynomial $|b(x_{g}y_{g}+\omega^{\prime})|$, whose irregularity is evaluated as
$\operatorname{irr}(|b(x_{g}y_{g}+\omega^{\prime})|)=\max\{\operatorname{irr}(|bx_{g}y_{g}|),\operatorname{irr}(|bx_{i}y_{i}|)\}=\operatorname{irr}(|bx_{g}y_{g}|)=\operatorname{irr}(|by_{g}x_{g}|)-1.$
This completes the proof.∎
Corollary 5.5.
For $a\in|T(H)|$, rewriting of the polynomial $a$ by $\rho$ terminates if and only if $\operatorname{irr}(a)<+\infty$.∎
Lemma 5.6.
Rewriting by $\rho$ is confluent: if $a^{\prime}_{1}$ and $a^{\prime}_{2}$ are obtained from $a\in|T(H)|$ by applying $\rho$ several times, there exists $a^{\prime\prime}\in|T(H)|$ by also applying $\rho$ several times that can be obtained either from $a^{\prime}_{1}$ or $a^{\prime}_{2}$.
Proof. It suffices to show where $a^{\prime}_{1}\neq a^{\prime}_{2}$ are obtained by only one application of $\rho$. By the assumption, there are two ways of applying $\rho$ to $a$: one leads to $a^{\prime}_{1}$ and the other to $a^{\prime}_{2}$. In this case, we can find distinct sequences of $y_{g}x_{g}$ in $a$ (either in the same monomial or different monomials) since it is impossible to share a part of sequence $y_{g}x_{g}$ without sharing the entirety of $y_{g}x_{g}$. Therefore, we can apply $\rho$ to the other sequence $y_{g}x_{g}$ in $a^{\prime}_{1}$ and $a^{\prime}_{2}$ respectively to obtain a common polynomial $a^{\prime\prime}$.∎
Corollary 5.7.
If the rewriting terminates, it leads to a unique polynomial. For $a\in|T(H)|$ with $\operatorname{irr}(a)$, we denote the thusly obtained element by $\rho(a)$.∎
Let us construct the basis of $|T(H)_{\omega}|$. From now on, the word “relation” refers to an expression of the form $\sum_{\lambda}c_{\lambda}w_{\lambda}\in|\langle\omega\rangle|$ where $c_{\lambda}\in\mathbb{K}$ and $w_{\lambda}\in C$. First of all, since the two-sided ideal $\langle\omega\rangle$ of $T(H)$ is generated by the sole element $\omega$, all the relations amongst elements of $C$ is given as a linear combination of the relation $|b\omega|=|b(x_{g}y_{g}+\omega^{\prime})|-|by_{g}x_{g}|\in|\langle\omega\rangle|$ for $b\in T(H)$, which come from the rule $\rho$ as a linear combinations of $w-a\in|\langle\omega\rangle|$ where $w\in C$ and $a$ is a polynomial obtained by applying $\rho$ to $w$.
If the rewriting terminates, we have a unique expression by the corollary above, so we have the normal form. In the case of $\operatorname{irr}(w)=+\infty$ for $w\in C$, however, several application of $\rho$ to $w$ can bring back to $w$ itself. For example, applying $\rho$ to the word $w=|y_{g}x_{g}x_{g}|$, we obtain
$a=|(x_{g}y_{g}+\omega^{\prime})x_{g}|=w+\sum_{1\leq i<g}|[x_{i},y_{i}]x_{g}|.$
Since $a$ and $w$ give the same element in $|T(H)_{\omega}|$, this yields a relation $\sum_{1\leq i<g}|[x_{i},y_{i}]x_{g}|\in|\langle\omega\rangle|$, which is a consequence of the rule $\rho$ but does not contain the sequence $y_{g}x_{g}$ in its expression. We deal with this type of relations from now on.
We need some preparation.
Definition 5.8.
Let $\mathcal{R}$ be a directed graph defined by the following:
  • Each vertex is labelled by $w\in C$ with $\operatorname{irr}(w)=+\infty$, which is a monomial consisting of $s$ $x_{g}$’s and $t$ $y_{g}$’s for some $s,t\geq 1$.
  • For one application of $\rho$ to $w$ with the result $w^{\prime}+a$, where $w^{\prime}\in C$ with $\operatorname{irr}(w)=+\infty$ and $a\in|T(H)|$ with $\operatorname{irr}(a)<+\infty$, we draw one directed edge from $w$ to $w^{\prime}$.
For an edge $e\colon w\mapsto w^{\prime}+a$, we define the holonomy $\mathrm{hol}(e)$ along the edge (with respect to the fixed gauges $C$) to be $\rho(a)$, which is well-defined by Corollary 5.7.
For $s,t\geq 1$, $\mathcal{R}^{(s,t)}$ denotes the full subgraph of $\mathcal{R}$ whose vertex labels are comprised of $s$ $x_{g}$’s and $t$ $y_{g}$’s. We have a decomposition $\mathcal{R}=\bigsqcup_{s,t\geq 1}\mathcal{R}^{(s,t)}$ as a directed graph.
A directed path is a (possibly empty) sequence of composable directed edges, and a loop is a directed path whose starting point coincides with the endpoint, which we call the base point of the loop. A free loop is an equivalence class of based loops generated by cyclic permutations of edges. The holonomy along a directed path is defined as the sum of the holonomies of edges comprising that path. For a pair of directed paths $w\mapsto w^{\prime}+a$ and $w\mapsto w^{\prime}+a^{\prime}$ in $\mathcal{R}$, we obtain a relation $a-a^{\prime}\in|\langle\omega\rangle|$. This is equivalent to the relation $\rho(a)-\rho(a^{\prime})\in|\langle\omega\rangle|$ between their holonomies since we already know $\rho(a)-a\in|\langle\omega\rangle|$ for any $a$ with $\operatorname{irr}(a)<+\infty$. All the relations are linearly generated by these, as mentioned above.
Lemma 5.9.
For two vertices $w,w^{\prime}$ in the same subgraph $\mathcal{R}^{(s,t)}$, there is a directed path from $w$ to $w^{\prime}$.
Proof. One application of $\rho$ to $w$ swaps one $y_{g}$ in $w$ with the $x_{g}$ immediately after this $y_{g}$. By applying $\rho$ many times, we can find a directed path from $w$ to $|x_{g}^{s}y_{g}^{t}|$. Now we can re-distribute $y_{g}$’s by applying $\rho$ many more times to reach the vertex $w^{\prime}$.∎
Lemma 5.10.
All the relations in $C$ are linearly generated by the relations obtained from loops in $\mathcal{R}$.
Proof. Given a pair of directed paths $w\mapsto w^{\prime}+a$ and $w\mapsto w^{\prime}+a^{\prime}$ considered above, we take a directed path $w^{\prime}\mapsto w+a^{\prime\prime}$, whose existence is guaranteed by the lemma above. Then, we obtain a pair of composite rewritings
$(w\mapsto w^{\prime}+a\mapsto w+a+a^{\prime\prime}\;,\;w\mapsto w^{\prime}+a^{\prime}\mapsto w+a^{\prime}+a^{\prime\prime}).$
This results in the relation $(a+a^{\prime\prime})-(a^{\prime}+a^{\prime\prime})\in|\langle\omega\rangle|$, which is exactly the relation $a-a^{\prime}\in|\langle\omega\rangle|$ obtained above. On the other hand, denoting the constant loop at $w$ by $\mathsf{c}_{w}$, the pairs of loops
$(w\mapsto w^{\prime}+a\mapsto w+a+a^{\prime\prime},\mathsf{c}_{w})\;\textrm{ and }\;(w\mapsto w^{\prime}+a^{\prime}\mapsto w+a+a^{\prime\prime},\mathsf{c}_{w})$
yield the relations $a+a^{\prime\prime},a^{\prime}+a^{\prime\prime}\in|\langle\omega\rangle|$, from which the relation $a-a^{\prime}\in|\langle\omega\rangle|$ follows. Therefore, relations coming from loops in $\mathcal{R}$ are enough to deduce all the relations. ∎
Lemma 5.11.
The holonomy along a loop does not depend on the base point.
Proof. For a loop traversing $w_{0},w_{1},\dotsc,w_{r}=w_{0}$ in this order with base point $w_{i}$, its holonomy is given by the sum $\sum_{0\leq i<r}a_{i}$, where $a_{i}$ is the holonomy along the edge from $w_{i}$ to $w_{i+1}$. This is clearly independent of the base point $w_{i}$.∎
Definition 5.12.
Let ~$\tilde{\mathcal{R}}^{(s,t)}$ be the directed graph defined by the following:
  • Each vertex is labelled by a cyclic sequence $\tilde{w}$ of letters $x_{g}^{[i]}$ and $y_{g}^{[j]}$ for $i\in\mathbb{Z}/s\mathbb{Z}$ and $j\in\mathbb{Z}/t\mathbb{Z}$ such that they appear exactly once while respecting the cyclic orders $(x_{g}^{[1]},\dotsc,x_{g}^{[s]})$ and $(y_{g}^{[1]},\dotsc,y_{g}^{[t]})$. The letters $x_{g}^{[i]}$ and $y_{g}^{[j]}$ are distinguished copies of $x_{g}$ and $y_{g}$, respectively.
  • For one application of $\rho$ to a cyclic sequence $\tilde{w}$ with the result $\tilde{w}^{\prime}+a$, where $w^{\prime}\in C$ with $\operatorname{irr}(\tilde{w}^{\prime})=+\infty$ and $a\in|T(H)|$ with $\operatorname{irr}(a)<+\infty$, we draw one edge from $\tilde{w}$ to $\tilde{w}^{\prime}$.
The natural quotient map ~$\tilde{\mathcal{R}}^{(s,t)}\twoheadrightarrow\mathcal{R}^{(s,t)}$ is induced by the re-labelling action of $\mathbb{Z}/s\mathbb{Z}\times\mathbb{Z}/t\mathbb{Z}$, which sends $x_{g}^{[i]}$ and $y_{g}^{[j]}$ appearing in labels to $x_{g}$ and $y_{g}$ respectively.
For a loop $\ell\colon w\mapsto w+a$ in $\mathcal{R}^{(s,t)}$, its $st$-fold power $\ell^{st}$ lifts to a loop in ~$\tilde{\mathcal{R}}^{(s,t)}$ and results in a rewriting $\tilde{w}\mapsto\tilde{w}+st\cdot a$, which gives a relation $st\cdot a\in|\langle\omega\rangle|$. Since we have assumed $\mathrm{char}(\mathbb{K})=0$, this is equlvalent to $a\in|\langle\omega\rangle|$. For this reason, it is enough to only consider loops in ~$\tilde{\mathcal{R}}^{(s,t)}$.
Lemma 5.13.
Let $w=|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(y_{g}^{[j]}x_{g}^{[i_{j}]}\cdots x_{g}^{[i_{j+1}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|$ be a vertex of ~$\tilde{\mathcal{R}}^{(s,t)}$. Then, we have a bijection
{edges in ~($\displaystyle\{\textrm{edges in }\tilde{\mathcal{R}}^{(s,t)}\textrm{ starting from }w\}\leftrightarrow\{j\in\mathbb{Z}/t\mathbb{Z}:i_{j}<i_{j+1}\}.$
Proof. An edge in ~$\tilde{\mathcal{R}}^{(s,t)}$ starting from $w$ can be written as the rewriting
$\displaystyle m_{j}$ $\displaystyle\colon|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(y_{g}^{[j]}x_{g}^{[i_{j}]}\cdots x_{g}^{[i_{j+1}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|$
$\displaystyle\quad\mapsto|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(x_{g}^{[i_{j}]}y_{g}^{[j]}x_{g}^{[i_{j}+1]}\cdots x_{g}^{[i_{j+1}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|+(\textrm{terms with }\operatorname{irr}<+\infty)\,.$
for some $j\in\mathbb{Z}/t\mathbb{Z}$. This moves $x_{g}^{[i_{j}]}$ to the left side of $y_{g}^{[j]}$ without changing other parts, and such a rewriting is possible only when $i_{j}<i_{j+1}$. The correspondence $m_{j}\leftrightarrow j$ yields the bijection we want.∎
As a consequence of the above lemma, any directed path in ~$\tilde{\mathcal{R}}^{(s,t)}$ can be expressed by a sequence $m_{j_{1}}\cdots m_{j_{r}}$, read from left to right. At this point, it should be convenient to introduce a diagrammatical explanation for the moves $m_{j}$. As in Figure 2, consider a circle whose arc is separated in $t$ parts by labelled partitions and $s$ beads are distributed over the segments. Each partition corresponds to one $y_{g}^{[j]}$, and each bead correspond to one $x_{g}^{[i]}$. One configuration of beads exactly specifies a unique vertex in ~$\tilde{\mathcal{R}}^{(s,t)}$. The rewriting $m_{j}$ removes the $i_{j}$-th bead, which is immediately right to the $j$-th partition, and adds to immediately left of the partition (Figure 2).
Figure 1: The configuration of beads corresponding to the vertex $|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|$.
Figure 2: The move $m_{j}$.
Lemma 5.14.
Two directed paths $m_{i}m_{j}$ and $m_{j}m_{i}$ have the same endpoints, and we have $\mathrm{hol}(m_{i}m_{j})=\mathrm{hol}(m_{j}m_{i})$.
Proof. Let $\tilde{w}=|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|$ for some $p,q$ and $b,b^{\prime}\in T(H)$. Then, the rewriting of $w^{\prime}$ along $m_{i}m_{j}$ reads
$\displaystyle|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|$ $\displaystyle\xmapsto{m_{i}}|b(x_{g}^{[p]}y_{g}^{[i]})b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|+|b\omega^{\prime}b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|$
$\displaystyle\xmapsto{m_{j}}|b(x_{g}^{[p]}y_{g}^{[i]})b^{\prime}(x_{g}^{[q]}y_{g}^{[j]})|+|b(x_{g}^{[p]}y_{g}^{[i]})b^{\prime}\omega^{\prime}|+|b\omega^{\prime}b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|,$
whose holonomy is
$\rho\big{(}|b(x_{g}y_{g})b^{\prime}\omega^{\prime}|+|b\omega^{\prime}b^{\prime}(y_{g}x_{g})|\big{)}=\rho\big{(}|b(x_{g}y_{g})b^{\prime}\omega^{\prime}|+|b\omega^{\prime}b^{\prime}(x_{g}y_{g}+\omega^{\prime})|\big{)}\,.$
On the other hand, the rewriting of $w^{\prime}$ along $m_{j}m_{i}$ reads
$\displaystyle|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}(y_{g}^{[j]}x_{g}^{[q]})|$ $\displaystyle\xmapsto{m_{j}}|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}(x_{g}^{[q]}y_{g}^{[j]})|+|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}\omega^{\prime})|$
$\displaystyle\xmapsto{m_{i}}|b(x_{g}^{[p]}y_{g}^{[i]})b^{\prime}(x_{g}^{[q]}y_{g}^{[j]})|+|b\omega^{\prime}b^{\prime}(x_{g}^{[q]}y_{g}^{[j]})|+|b(y_{g}^{[i]}x_{g}^{[p]})b^{\prime}\omega^{\prime}|$
whose holonomy is
$\rho\big{(}|b\omega^{\prime}b^{\prime}(x_{g}y_{g})|+|b(y_{g}x_{g})b^{\prime}\omega^{\prime}|\big{)}=\rho\big{(}|b\omega^{\prime}b^{\prime}(x_{g}y_{g})|+|b(x_{g}y_{g}+\omega^{\prime})b^{\prime}\omega^{\prime}|\big{)}\,.$
The endpoints and their holonomies are equal, as is claimed.∎
From now on, we study the set of directed paths in modulo the relation $m_{i}m_{j}=m_{j}m_{i}$.
Definition 5.15.
For a vertex $w$ of ~$\tilde{\mathcal{R}}^{(s,t)}$, we set $P^{(s,t)}_{w}=\{\textrm{directed paths in }\tilde{\mathcal{R}}^{(s,t)}\textrm{ starting from }w\}$. We say a directed path $p_{1}$ is a deformation of $p_{2}$ if $p_{1}$ and $p_{2}$ give the same element in $P^{(s,t)}_{w}/\langle m_{i}m_{j}=m_{j}m_{i}\rangle$, which is the quotient by the relations multiplicatively generated by $m_{i}m_{j}=m_{j}m_{i}$ for all $i,j$, whenever defined.
By the lemma above, the holonomy map is well-defined on the set $P^{(s,t)}_{w}/\langle m_{i}m_{j}=m_{j}m_{i}\rangle$.
Definition 5.16.
We define the counting number map by
$\displaystyle\mathbf{c}\colon P^{(s,t)}_{w}\to(\mathbb{Z}_{\geq 0})^{t}\colon p\mapsto(\mathbf{c}(p,j))_{j\in\mathbb{Z}/t\mathbb{Z}}$
where $\mathbf{c}(p,j)=(\textrm{the number of beads jumped over }y_{g}^{[j]}\textrm{ along }p)=(\textrm{the number of }m_{j}\textrm{'s in }p)$.
Lemma 5.17.
The counting number map induces an injection $P^{(s,t)}_{w}/\langle m_{i}m_{j}=m_{j}m_{i}\rangle\to(\mathbb{Z}_{\geq 0})^{t}$.
Proof. Let $w=|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|$. Suppose that two directed paths $p$ and $p^{\prime}$ have the same counting number: $\mathbf{c}(p)=\mathbf{c}(p^{\prime})$. Put $\nu=\sum_{1\leq j\leq t}\mathbf{c}(p,j)$, which is the total number of $m_{j}$’s appearing in each of $p$ and $p^{\prime}$. We can write as $p=m_{j_{1}}\cdots m_{j_{\nu}}$ and $p^{\prime}=m_{j^{\prime}_{1}}\cdots m_{j^{\prime}_{\nu}}$. By the assumption $\mathbf{c}(p)=\mathbf{c}(p^{\prime})$, $p$ and $p^{\prime}$ has the same number of $m_{j}$’s for each $j$.
We prove by induction on $\nu$. If $\nu=0$, they are constant loops at $w$, so they are equal. Now suppose $\nu\geq 1$. Let $\alpha\in\{1,\dotsc,\nu\}$ such that $m_{j^{\prime}_{\alpha}}$ is the leftmost appearence of $m_{j_{1}}$ in $p^{\prime}$. Since $m_{j_{1}}$ is the first edge of $p$, we have $i_{j_{1}}<i_{j_{1}+1}$, which says that there is at least one bead between $j_{1}$-th and $(j_{1}+1)$-th partitions. For $1\leq\beta<\alpha$, we have $j^{\prime}_{\beta}\neq j_{1}$ so the move $m_{j^{\prime}_{\beta}}$ does not reduce the number of beads between $j_{1}$-th and $(j_{1}+1)$-th partitions. Therefore, we have $m_{j^{\prime}_{\beta}}m_{j^{\prime}_{\alpha}}=m_{j^{\prime}_{\alpha}}m_{j^{\prime}_{\beta}}$ for all $1\leq\beta<\alpha$; hence we can assume $j^{\prime}_{1}=j_{1}$. Now we can use the induction hypothesis to $m_{j_{2}}\cdots m_{j_{\nu}}$ and $m_{j^{\prime}_{2}}\cdots m_{j^{\prime}_{\nu}}$ to conclude that they are equal modulo the relation $m_{i}m_{j}=m_{j}m_{i}$. This shows $p=p^{\prime}$ in $P_{w}^{(s,t)}$.∎
We have a preferred base point $*_{s,t}=|y_{g}^{[1]}\cdots y_{g}^{[t]}x_{g}^{[1]}\cdots x_{g}^{[s]}|$ in ~$\tilde{\mathcal{R}}^{(s,t)}$ together with a based loop $m_{t}^{s}m_{t-1}^{s}\cdots m_{1}^{s}$, which first passes the edge specified by $t\in\mathbb{Z}/t\mathbb{Z}$ exactly $s$ times, then $(t-1)\in\mathbb{Z}/t\mathbb{Z}$ exactly $s$ times, and so on, starting at the above base point.
Figure 3: The collection move and the re-distribution move.
Lemma 5.18.
For any non-constant loop $\ell$ in ~$\tilde{\mathcal{R}}^{(s,t)}$ represented by a composition of $m_{j}$’s, there is a sequence of deformations and cyclic permutations of $m_{j}$’s which converts $\ell$ to $(m_{t}^{s}m_{t-1}^{s}\cdots m_{1}^{s})^{n}$ for some $n\geq 1$.
Proof. Let $w=|(y_{g}^{[1]}x_{g}^{[i_{1}]}\cdots x_{g}^{[i_{2}-1]})\cdots(y_{g}^{[t]}x_{g}^{[i_{t}]}\cdots x_{g}^{[i_{1}-1]})|$ be a base point of $\ell$. Since this is non-constant, at least one bead is moved by $\ell$. However, since $\ell$ is a loop, that one bead has to go around the circle to get back to the original position; this forces every bead, especially $x_{g}^{[1]}$, to go around the circle. We have $\mathbf{c}(\ell)=(ns,ns,\dotsc,ns)$ for some $n\in\mathbb{Z}_{>0}$.
Now consider the collection move and the re-distribution move depicted in Figure 3. Let $j$ be such that $i_{j}\leq 1<i_{j+1}$. The collection move is given by $m_{w}^{\mathrm{col}}=m_{j}^{r_{w,j}}m_{j-1}^{r_{w,j-1}}\cdots m_{j+1}^{r_{w,j+1}}$ where $r_{w,k}=s+1-i_{k}\in\{0,\dotsc,s-1\}$. This move gathers all the beads in $j$-th segment satisfying $i_{j}\leq 1<i_{j+1}$ by pushing the red line in the figure around the circle so that it ends at the vertex $|y_{g}^{[j+1]}\cdots y_{g}^{[j]}x_{g}^{[1]}\cdots x_{g}^{[s]}|$. The re-distribution move is given by $m_{w}^{\mathrm{red}}=m_{j}^{s-r_{w,j}}m_{j-1}^{s-r_{w,j-1}}\cdots m_{j+1}^{s-r_{w,j+1}}$. This move sends the beads in bulk along the circle until they reach the original configuration $w$. Then, the composition $\ell^{\prime}=(m_{w}^{\mathrm{col}}m_{w}^{\mathrm{red}})^{n}$ is a loop, and the counting number is $\mathbf{c}(\ell^{\prime})=(ns,\dotsc,ns)=\mathbf{c}(\ell)$. By Lemma 5.17, $\ell$ is a deformation of $\ell^{\prime}$.
Put ′′$\ell^{\prime\prime}=(m_{w}^{\mathrm{red}}m_{w}^{\mathrm{col}})^{n}$, which is a cyclic permutation of $\ell^{\prime}$ based at $|y_{g}^{[j+1]}\cdots y_{g}^{[j]}x_{g}^{[1]}\cdots x_{g}^{[s]}|$. On the other hand, $(m_{j}^{s}m_{j-1}^{s}\cdots m_{j+1}^{s})^{n}$ also has the same base point and the counting number. Again by Lemma 5.17, $\ell^{\prime\prime}$ is a deformation of $(m_{j}^{s}m_{j-1}^{s}\cdots m_{j+1}^{s})^{n}$. Finally, this is a cyclic permutation of the one in the claim. ∎
Corollary 5.19.
For any loop $\ell$ in ~$\tilde{\mathcal{R}}^{(s,t)}$, we have $\mathrm{hol}(\ell)=n\cdot\mathrm{hol}(m_{t}^{s}m_{t-1}^{s}\cdots m_{1}^{s})$ for some $n\in\mathbb{Z}$.∎
Computation 5.20.
We compute $h=\mathrm{hol}(m_{t}^{s}m_{t-1}^{s}\cdots m_{1}^{s})$. The path $m_{j}^{s}$ projects to the same loop $\ell_{(s,t)}$ in $\mathcal{R}^{(s,t)}$ independent of $j\in\mathbb{Z}/t\mathbb{Z}$, so we have $h=t\cdot\mathrm{hol}(\ell_{(s,t)})$. Therefore, it is enough to compute $\mathrm{hol}(m_{t}^{s})$. We have
$\displaystyle|y_{g}^{[1]}\cdots y_{g}^{[t]}x_{g}^{[1]}\cdots x_{g}^{[s]}|$ $\displaystyle\xmapsto{m_{t}}|y_{g}^{[1]}\cdots y_{g}^{[t-1]}(x_{g}^{[1]}y_{g}^{[t]}+\omega^{\prime})x_{g}^{[2]}\cdots x_{g}^{[s]}|$
$\displaystyle\xmapsto{m_{t}}\cdots$
$\displaystyle\xmapsto{m_{t}}|y_{g}^{[1]}\cdots y_{g}^{[t-1]}x_{g}^{[1]}\cdots x_{g}^{[s]}y_{g}^{[t]}+\sum_{1\leq i\leq s}y_{g}^{[1]}\cdots y_{g}^{[t-1]}x_{g}^{[1]}\cdots x_{g}^{[i-1]}\omega^{\prime}x_{g}^{[i+1]}\cdots x_{g}^{[s]}|.$
Put $r_{s,t}=\sum_{1\leq i\leq s}x_{g}^{s-i}y_{g}^{t-1}x_{g}^{i-1}\in T(H)$ for $s,t\geq 1$. Since $\operatorname{irr}(|r_{s,t}\omega^{\prime}|)<+\infty$ and $\omega^{\prime}$ does not contain any $x_{g}$’s and $y_{g}$’s, we can reduce $\operatorname{irr}(|r_{s,t}\omega^{\prime}|)$ to 00 0 by applying $\rho$ to $r_{s,t}$ inside $|r_{s,t}\omega^{\prime}|$. We put $|r^{\prime}_{s,t}\omega^{\prime}|=\rho(|r_{s,t}\omega^{\prime}|)$; we have $h=t\cdot|r^{\prime}_{s,t}\omega^{\prime}|$. In partiular, if $s=t=1$, we have $r_{s,t}=r^{\prime}_{s,t}=1$, which yields $h=|\omega^{\prime}|=0$. This gives the trivial relation $0\in|\langle\omega\rangle|$, so the case $s=t=1$ can be discarded.
We write $r^{\prime}_{s,t}$ in the form $r^{\prime}_{s,t}=sx_{g}^{s-1}y_{g}^{t-1}+b_{s,t}$. Then, $b_{s,t}\in T(H)$ has irregularity 00 0 and the number of $x_{g}$’s and $y_{g}$’s in each term in $b_{s,t}$ is less than $(s-1)$ and $(t-1)$, respectively.
With this being seen, we define a secondary rewriting rule of monomials (not of a sequence in a word!) for $g\geq 2$ and $s,t\geq 1$, $s+t\geq 3$, by
$|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|\xmapsto{\rho_{2}}|s^{-1}r^{\prime}_{s,t}(x_{g-1}y_{g-1}+\omega^{\prime\prime})-s^{-1}b_{s,t}y_{g-1}x_{g-1}|\,,$
where $\omega^{\prime\prime}=\sum_{1\leq i<g-1}[x_{i},y_{i}]$. Note that the terms on the right-hand side have irregularity 00 0, and $s^{-1}$ is well-defined since we assumed $\mathrm{char}(\mathbb{K})=0$. This rewriting eliminates the monomials $|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|$ since these do not appear in the right-hand side of $\rho_{2}$ by the condition $s+t\geq 3$. It always terminates since a polynomial has a finite number of terms.
Remark 5.21.
If $s=t=1$, the definition of $\rho_{2}$ reads $|y_{g-1}x_{g-1}|\xmapsto{\rho_{2}}|x_{g-1}y_{g-1}|$, which rewrites nothing at all. Therefore, the monomial $|y_{g-1}x_{g-1}|$ cannot be eliminated.
Now, we state the main result of this section.
Theorem 5.22.
Define the subset $X$, $Y$ of $C$ by
$\displaystyle X$ =$\displaystyle=\{w\in C:\operatorname{irr}(w)=0,w\neq|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|\textrm{ for any }s,t\geq 1\textrm{ with }s+t\geq 3\},\textrm{ and}$
$\displaystyle Y$ =$\displaystyle=\{|x_{g}^{s}y_{g}^{t}|\in C:s,t\geq 1\}.$
Then, the image of $X\sqcup Y$ in $|T(H)_{\omega}|$ is a basis for any $g\geq 1$.
We need another kind of graph for the proof.
Definition 5.23.
  • A skew-weighted graph is a triple $\mathcal{G}=(V(\mathcal{G}),E(\mathcal{G}),\mathsf{wt})$ where $(V(\mathcal{G}),E(\mathcal{G}))$ is a finite directed graph and $\mathsf{wt}$ is a weight map $\mathsf{wt}\colon\mathrm{HE}(\mathcal{G})\to\mathbb{K}$. Here we denote by $\mathrm{HE}(\mathcal{G})=\bigsqcup_{w\in V(\mathcal{G})}\mathrm{HE}(w)$ the set of half-edges. We require every vertex of $\mathcal{G}$ to be at least univalent, and $\mathsf{wt}(h)=-\mathsf{wt}(\bar{h})$ for each pair of half edges $(h,\bar{h})$ comprising an edge. Graphs may have self-loops (i.e. an edge whose half-edges are both located at the same vertex) and multiple edges. For $h\in\mathrm{HE}(\mathcal{G})$, which is a part of a unique edge $\lambda$, we set $\varepsilon_{h}=+1$ if $h$ is the first half of the directed edge $\lambda$ and $-1$ otherwise.
  • A cycle decomposition of a skew-weighted graph $\mathcal{G}$ is a finite family $(\mathcal{G}_{z})_{z\in Z}$ of skew-weighted bivalent subgraphs of $\mathcal{G}$, each with a choice of global orientation (out of possible two of a circle). The weight map $\mathsf{wt}_{z}$ on $\mathcal{G}_{z}$ is not necessarily a restriction of $\mathsf{wt}$. Put $\mathrm{HE}(\mathcal{G}_{z})=\{h^{z}_{1},\bar{h}^{z}_{1},\dotsc,h^{z}_{r_{z}},\bar{h}^{z}_{r_{z}}\}$ in this cyclic order with respect to the chosen global orientation, where $h^{z}_{i}$ and $\bar{h}^{z}_{i}$ is contained in the same edge $\lambda^{z}_{i}$. We require that $c_{z}:=\mathsf{wt}_{z}(h^{z}_{i})$ is independent of $i\in\{1,\dotsc,r_{z}\}$, $\mathcal{G}=\bigcup_{z\in Z}\mathcal{G}_{z}$ as a usual graph and
    $\displaystyle\mathsf{wt}(h)=\sum_{z\in Z:\,h\in\mathrm{HE}(\mathcal{G}_{z})}\mathsf{wt}_{z}(h)$ (2)
    for each $h\in\mathrm{HE}(\mathcal{G})$.
Lemma 5.24.
Let $\mathcal{G}$ be a skew-weighted graph such that $\mathsf{wt}(h)\neq 0$ for all $h\in\mathrm{HE}(\mathcal{G})$ and
$\displaystyle\sum_{h\in\mathrm{HE}(w)}\mathsf{wt}(h)=0$ (3)
for all $w\in V(\mathcal{G})$. Then, there exists a cyclic decomposition of $\mathcal{G}$.
Proof. If $\mathcal{G}$ has a univalent vertex $w$, there is a unique $h\in\mathrm{HE}(w)$ and we have $\mathsf{wt}(h)=0$, which contradicts the assumption. Therefore, $\mathcal{G}$ has no univalent vertices.
We proceed by induction on $n=\#E(\mathcal{G})$. If $n=0$, there are no edges. Since we assumed that every vertex is at least univalent, the graph itself is empty. Let $n\geq 1$ and take a bivalent subgraph $\mathcal{G}_{1}$ in $\mathcal{G}$. If there are no such subgraphs, $\mathcal{G}$ is necessarily a tree, which is impossible since $\mathcal{G}$ has no univalent vertex. Choose a global orientation and put $\mathrm{HE}(\mathcal{G}_{1})=\{h_{1},\bar{h}_{1},\dotsc,h_{r},\bar{h}_{r}\}$ in this cyclic order. Put $c=\mathsf{wt}(h_{1})$. We define a weight map on $\mathcal{G}_{1}$ by $\mathsf{wt}_{1}(h_{i})=c$. Then $(\mathcal{G}_{1},\mathsf{wt}_{1})$ is a skew-weighted graph. We define another weight map on $\mathcal{G}$ by
$\mathsf{wt}_{2}(h)=\left\{\begin{aligned} &\mathsf{wt}(h)-\mathsf{wt}_{1}(h)&&\;\textrm{ if }h\in\mathrm{HE}(\mathcal{G}^{\prime}),\\ &\mathsf{wt}(h)&&\;\textrm{ otherwise.}\end{aligned}\right.$
Let $(\mathcal{G}_{2},\mathsf{wt}_{2})$ be the skew-weighted subgraph of $\mathcal{G}$ specified by $E(\mathcal{G}_{2})=\{\lambda=(h,\bar{h})\in E(\mathcal{G}):\mathsf{wt}_{2}(h)\neq 0\}$. Since $\mathsf{wt}_{2}(h_{1})=0$, we have #$\#E(\mathcal{G}_{2})<n$. Since $\mathcal{G}_{2}$ also satisfies the condition (3), we can take a cyclic decomposition $(\mathcal{G}_{2,z})_{z\in Z_{2}}$ of $\mathcal{G}_{2}$ by the induction hypothesis. It is readily seen that $(\mathcal{G}_{2,z})_{z\in Z_{2}}\cup\{\mathcal{G}_{1}\}$ is a cyclic decomposition. ∎
Proof of Theorem 5.22. The case of $g=1$ is straightforward, so we assume $g\geq 2$. We first show the linear independence of the image of $X$ in $|T(H)_{\omega}|$, which is equivalent to $\mathrm{Span}_{\mathbb{K}}(X)\cap|\langle\omega\rangle|=0$ in $|T(H)|$. Take $a\in\mathrm{Span}_{\mathbb{K}}(X)\cap|\langle\omega\rangle|$, which can be expressed as $a=\sum_{\lambda\in\Lambda}c_{\lambda}(w_{\lambda}-a_{\lambda})$ where $\Lambda$ is a finite set, $c_{\lambda}\in\mathbb{K}\setminus\{0\}$, $w_{\lambda}\in C$ with arbitrary irregularity (including +$\infty$+ ∞) and $a_{\lambda}\in|T(H)|$ is obtained by applying $\rho$ more than one times to $w_{\lambda}$. Set $\Lambda_{n}=\{\lambda\in\Lambda:\operatorname{irr}(w_{\lambda})=n\}$; we have $\operatorname{irr}(w_{\lambda})>\operatorname{irr}(a_{\lambda})$ for $\lambda\in\Lambda_{n}$ with $n<+\infty$. For $\lambda\in\Lambda_{+\infty}$, we put $a_{\lambda}=w^{\prime}_{\lambda}+a^{\prime}_{\lambda}$ where $w^{\prime}_{\lambda}\in C$ with $\operatorname{irr}(w^{\prime}_{\lambda})=+\infty$ and $\operatorname{irr}(a^{\prime}_{\lambda})<+\infty$, and we have an associated path $p_{\lambda}:w_{\lambda}\mapsto w^{\prime}_{\lambda}$ in $\mathcal{R}$ by remembering how we obtained $a_{\lambda}$ from $w_{\lambda}$. Setting $C_{n}=\{w\in C:\operatorname{irr}(w)=n\}$ for $n\in\mathbb{Z}_{\geq 0}\cup\{+\infty\}$, we have a projection
$\mathrm{proj}_{+\infty}\colon|T(H)|\to\mathrm{Span}_{\mathbb{K}}(C_{+\infty})$
induced by the decomposition $C=C_{+\infty}\sqcup(C\setminus C_{+\infty})$. Then, since $\operatorname{irr}(a)=0$, we have
$\displaystyle 0=\mathrm{proj}_{+\infty}(a)$ $\displaystyle=\mathrm{proj}_{+\infty}\Big{(}\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}(w_{\lambda}-a_{\lambda})\Big{)}=\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}(w_{\lambda}-w^{\prime}_{\lambda}).$ (4)
Therefore, we have
$\displaystyle a=\sum_{\lambda\in\Lambda\setminus\Lambda_{+\infty}}c_{\lambda}(w_{\lambda}-a_{\lambda})+\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}(-a^{\prime}_{\lambda}).$
Since $a\in\mathrm{Span}_{\mathbb{K}}(X)$, we have $\operatorname{irr}(a)=0$ and
$\displaystyle a=\rho(a)$ $\displaystyle=\rho\Big{(}\sum_{\lambda\in\Lambda\setminus\Lambda_{+\infty}}c_{\lambda}(w_{\lambda}-a_{\lambda})+\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}(-a^{\prime}_{\lambda})\Big{)}$
$\displaystyle=\sum_{\lambda\in\Lambda\setminus\Lambda_{+\infty}}c_{\lambda}(\rho(w_{\lambda})-\rho(a_{\lambda}))-\rho\Big{(}\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}a^{\prime}_{\lambda}\Big{)}$
$\displaystyle=-\rho\Big{(}\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}a^{\prime}_{\lambda}\Big{)}.$
Here we used $\rho(w_{\lambda})=\rho(a_{\lambda})$ for $\lambda\in\Lambda\!\setminus\!\Lambda_{+\infty}$, which follows from Corollary 5.7.
Now, we construct a skew-weighted graph. Set $V(\mathcal{G})=\{w_{\lambda},w^{\prime}_{\lambda}:\lambda\in\Lambda_{+\infty}\}\subset C$ and $E(\mathcal{G})=\Lambda_{+\infty}$. Each $\lambda$ corresponds to an edge from $w_{\lambda}$ to $w^{\prime}_{\lambda}$. The weight map is defined by $\mathsf{wt}(h)=\varepsilon_{h}c_{\lambda}$ where $h$ is a part of a unique edge $\lambda$ from $w_{\lambda}$ and $w^{\prime}_{\lambda}$. The equality (4) is equivalent to
$\displaystyle\sum_{h\in\mathrm{HE}(w)}\mathsf{wt}(h)=0$
for each $w\in V(\mathcal{G})$. Since $\mathsf{wt}(h)\neq 0$ for all $h$ by the assumption, we can take a cycle decomposition $(\mathcal{G}_{z})_{z\in Z}$ of $\mathcal{G}$ by Lemma 5.24. The right-hand side of (2) is equal to
$\sum_{z\in Z:\,h=h_{i}^{z}}\mathsf{wt}_{z}(h)+\sum_{z\in Z:\,h=\bar{h}_{i}^{z}}\mathsf{wt}_{z}(h)=\sum_{z\in Z:\,h=h_{i}^{z}}c_{z}-\sum_{z\in Z:\,h=\bar{h}_{i}^{z}}c_{z},$
so we obtain
$\displaystyle\sum_{\lambda=(h,\bar{h})\in\Lambda_{+\infty}}c_{\lambda}(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})$
$\displaystyle=\sum_{\lambda=(h,\bar{h})\in\Lambda_{+\infty}}\varepsilon_{h}\mathsf{wt}(h)(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})$
$\displaystyle=\sum_{\lambda=(h,\bar{h})\in\Lambda_{+\infty}}\bigg{(}\sum_{z\in Z:\,h=h_{i}^{z}}\varepsilon_{h}c_{z}(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})-\sum_{z\in Z:\,h=\bar{h}_{i}^{z}}\varepsilon_{h}c_{z}(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})\bigg{)}$
$\displaystyle=\sum_{z\in Z}\bigg{(}\sum_{\begin{subarray}{c}\lambda=(h,\bar{h})\in\Lambda_{+\infty}\\ h=h_{i}^{z}\end{subarray}}\varepsilon_{h}c_{z}(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})-\sum_{\begin{subarray}{c}\lambda=(h,\bar{h})\in\Lambda_{+\infty}\\ h=\bar{h}_{i}^{z}\end{subarray}}\varepsilon_{h}c_{z}(w_{\lambda}-w^{\prime}_{\lambda}-a^{\prime}_{\lambda})\bigg{)}$
$\displaystyle=\sum_{z\in Z}c_{z}\bigg{(}\sum_{\begin{subarray}{c}\lambda=(h,\bar{h})\in\Lambda_{+\infty}\\ h=h_{i}^{z}\end{subarray}}\varepsilon_{h_{i}^{z}}(w_{\lambda^{z}_{i}}-w^{\prime}_{\lambda^{z}_{i}}-a^{\prime}_{\lambda^{z}_{i}})+\sum_{\begin{subarray}{c}\lambda=(h,\bar{h})\in\Lambda_{+\infty}\\ h=\bar{h}_{i}^{z}\end{subarray}}\varepsilon_{h_{i}^{z}}(w_{\lambda^{z}_{i}}-w^{\prime}_{\lambda^{z}_{i}}-a^{\prime}_{\lambda^{z}_{i}})\bigg{)}$
$\displaystyle=\sum_{z\in Z}c_{z}\sum_{1\leq i\leq r_{z}}\varepsilon_{h_{i}^{z}}(w_{\lambda^{z}_{i}}-w^{\prime}_{\lambda^{z}_{i}}-a^{\prime}_{\lambda^{z}_{i}}).$
We can choose the global orientation so that $\varepsilon_{h^{z}_{1}}=+1$. We set $a^{\prime}_{z}=\sum_{1\leq i\leq r_{z}}\varepsilon_{h_{i}^{z}}a^{\prime}_{\lambda^{z}_{i}}$ for $z\in Z$ so that we have
$\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}a^{\prime}_{\lambda}=\sum_{z\in Z}c_{z}a^{\prime}_{z}.$
From a bivalent subgraph $\mathcal{G}_{z}$, we construct a genuine loop $\ell_{z}$ in ~$\tilde{\mathcal{R}}^{(s,t)}$ inductively by the following. We start at $w_{\lambda^{z}_{1}}$ and traverse along the path $p_{\lambda^{z}_{1}}$ to reach $w^{\prime}_{\lambda^{z}_{1}}$. For $i\geq 2$, if the next path $p_{\lambda^{z}_{i}}$ is in the correct direction (namely, from $w_{\lambda^{z}_{i}}$ to $w^{\prime}_{\lambda^{z}_{i}}$, in which case $\varepsilon_{h_{i}^{z}}=+1$), traverse $p_{\lambda^{z}_{i}}$; if not, traverse along an arbitrary directed path $p^{\prime}_{\lambda^{z}_{i}}$ from $w_{\lambda^{z}_{i}}$ to $w^{\prime}_{\lambda^{z}_{i}}$, whose existence is guaranteed by Lemma 5.9. Then, the holonomy along $\ell_{z}$ is equal to a scalar multiple of $|r^{\prime}_{s,t}\omega^{\prime}|$ by Corollary 5.19 and Computation 5.20. For $\lambda^{z}_{i}$ such that $\varepsilon_{h^{z}_{i}}=-1$, we have a loop $p_{\lambda^{z}_{i}}p^{\prime}_{\lambda^{z}_{i}}$ in $\mathcal{R}$, whose holonomy is given by $\mathrm{hol}(p_{\lambda^{z}_{i}})+\mathrm{hol}(p^{\prime}_{\lambda^{z}_{i}})=q_{z,i}|r^{\prime}_{s_{z},t_{z}}\omega^{\prime}|$ for some $q_{z,i}\in\mathbb{K}$. Then, we have
$\displaystyle a^{\prime}_{z}-\mathrm{hol}(\ell_{z})$ $\displaystyle=\sum_{1\leq i\leq r_{z}}\varepsilon_{h^{z}_{i}}\mathrm{hol}(p_{\lambda^{z}_{i}})-\sum_{\begin{subarray}{c}1\leq i\leq r_{z}\\ \varepsilon_{h^{z}_{i}}=+1\end{subarray}}\mathrm{hol}(p_{\lambda^{z}_{i}})-\sum_{\begin{subarray}{c}1\leq i\leq r_{z}\\ \varepsilon_{h^{z}_{i}}=-1\end{subarray}}\mathrm{hol}(p^{\prime}_{\lambda^{z}_{i}})$
$\displaystyle=\sum_{\begin{subarray}{c}1\leq i\leq r_{z}\\ \varepsilon_{h^{z}_{i}}=-1\end{subarray}}\varepsilon_{h^{z}_{i}}\mathrm{hol}(p_{\lambda^{z}_{i}})-\mathrm{hol}(p^{\prime}_{\lambda^{z}_{i}})$
$\displaystyle=\sum_{\begin{subarray}{c}1\leq i\leq r_{z}\\ \varepsilon_{\lambda_{i}}=-1\end{subarray}}-q_{z,i}|r^{\prime}_{s_{z},t_{z}}\omega^{\prime}|\,.$
Therefore, we have $\rho(a^{\prime}_{z})=q_{z}|r^{\prime}_{s_{z},t_{z}}\omega^{\prime}|$ for some $q_{z}\in\mathbb{K}$. Hence, we obtain
$a=-\rho\Big{(}\sum_{\lambda\in\Lambda_{+\infty}}c_{\lambda}a^{\prime}_{\lambda}\Big{)}=-\sum_{z\in Z}c_{z}\rho(a^{\prime}_{z})=-\sum_{z\in Z}c_{z}q_{z}|r^{\prime}_{s_{z},t_{z}}\omega^{\prime}|=-\sum_{s,t\geq 1}c_{s,t}|r^{\prime}_{s,t}\omega^{\prime}|=-\sum_{\begin{subarray}{c}s,t\geq 1\\ s+t\geq 3\end{subarray}}c_{s,t}|r^{\prime}_{s,t}\omega^{\prime}|$
for some $c_{s,t}\in\mathbb{K}$. We introduce another projection
$\mathrm{proj}^{\prime}\colon\mathrm{Span}_{\mathbb{K}}(C_{0})\to\mathrm{Span}_{\mathbb{K}}(C_{0}\setminus X)$
induced from the decomposition $C_{0}=X\sqcup(C_{0}\setminus X)$. Since $a\in\mathrm{Span}_{\mathbb{K}}(X)$, We have
$0=\mathrm{proj}^{\prime}(a)=\sum_{\begin{subarray}{c}s,t\geq 1\\ s+t\geq 3\end{subarray}}c_{s,t}\cdot s|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|.$
The monomials $|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|$ are distinct, so we obtain $c_{s,t}\cdot s=0$ for all $s,t\geq 1$ with $s+t\geq 3$, which is equivalent to $c_{s,t}=0$. Therefore, we obtain $a=0\,$; this shows the linear independency of the image of $X$ in $|T(H)_{\omega}|$.
Next, considering the $\mathbb{K}$-algebra map
$\phi\colon T(H)_{\omega}\to\mathbb{K}[x_{g},y_{g}]\colon x_{i},y_{i}\mapsto 0\,(i<g),\;x_{g}\mapsto x_{g},\;y_{g}\mapsto y_{g}\,,$
the composite
$\mathrm{Span}_{\mathbb{K}}(X\sqcup Y)\to|T(H)_{\omega}|\xrightarrow{|\phi|}\mathbb{K}[x_{g},y_{g}]\twoheadrightarrow\mathbb{K}[x_{g},y_{g}]/\langle x_{g}^{s},y_{g}^{t}\colon s,t\geq 0\rangle_{\mathbb{K}\textrm{-vect}}$
sends $X$ to 0 and maps $\mathrm{Span}_{\mathbb{K}}(Y)$ to the image isomorphically. Therefore, the image of $X\sqcup Y$ in $|T(H)_{\omega}|$ is linearly independent.
Finally, we show $\mathrm{Span}_{\mathbb{K}}(X\sqcup Y)$ surjects onto $|T(H)_{\omega}|$. For $a\in|T(H)|$, apply $\rho$ to the terms in $a$ with $\operatorname{irr}=+\infty$ so that they are contained in $\mathrm{Span}_{\mathbb{K}}(Y)$. Denote the (not necessarily unique) result by $a^{\prime}$. Next, apply $\rho$ to the terms in $a^{\prime}$ with $\operatorname{irr}<+\infty$ until their irregularity becomes zero. Denoting the result by $a^{\prime\prime}$, apply $\rho_{2}$ to all the terms in $a^{\prime\prime}$ which are scalar multiples of $|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|$ for $s+t\geq 3$ to eliminate them. The final result $a^{\prime\prime\prime}$ is now contained in $\mathrm{Span}_{\mathbb{K}}(X\sqcup Y)$, while $a^{\prime},a^{\prime},a^{\prime\prime}$ and $a^{\prime\prime\prime}$ are all equal modulo $|\langle\omega\rangle|$. ∎
6. The Closed Surface Case
In this section, we solve the formality problem for a closed surface $\Sigma=\Sigma_{g,0}$. The following is already known.
Theorem 6.1.
The completed Goldman–Turaev Lie bialgebra $|\widehat{\mathbb{K}\pi}/\mathbb{K}1|$ for a closed surface is formal.
Proof. This is a consequence of the formality on a surface with only one boundary component. Since a framing does not matter in this case, the set of formality isomorphisms is non-empty by Theorem 7.1 of [AKKN23] for any $g\geq 1$.∎
The theorem above states that the set of formality isomorphisms is a torsor over their pro-unipotent automorphism groups. At this point, describing the automorphism group suffices.
We are done with linear categories and return to the usual algebras. Let $\hat{L}(H)_{\omega}=\hat{L}(H)/\langle\omega\rangle$ be the quotient Lie algebra by the ideal generated by $\omega$, and $\hat{T}(H)_{\omega}=U\hat{L}(H)_{\omega}$, which is a quotient of $\hat{T}(H)$ by the (complete) two-sided ideal generated by $\omega$. Now, we take the standard resolution of $\mathbb{K}$ as a left $\hat{T}(H)_{\omega}$-module
$\displaystyle 0\to P_{1}\xrightarrow{\partial_{1}}$ $\displaystyle P_{0}\xrightarrow{\partial_{0}}\hat{T}(H)_{\omega}\xrightarrow{\mathrm{aug}}\mathbb{K}\to 0$
where $P_{0}=\hat{T}(H)_{\omega}\otimes H$, $P_{1}=\hat{T}(H)_{\omega}\otimes\mathbb{K}\omega$, and
$\partial_{1}(1\otimes\omega)=\sum_{i}x_{i}\otimes y_{i}-y_{i}\otimes x_{i}\,,\;\partial_{0}(a\otimes v)=av\,.$
We define connections on each piece:
$\displaystyle\nabla^{\prime}_{0,H}(1\otimes x_{i})=\nabla^{\prime}_{0,H}(1\otimes y_{i})=0\textrm{ on }P_{0}\textrm{ and }\nabla^{\prime}_{1,H}(1\otimes\omega)=0\textrm{ on }P_{1}\,.$
Put $\nabla^{\prime}_{\bullet,H}=\{\nabla^{\prime}_{i,H}\}_{i\geq 0}$; this comprises a flat homological connection defined in [Tan24b], Section 7.
The following is the main result of this paper.
Theorem 6.2.
The pro-unipotent part of the automorphism group of the associated graded of the Goldman–Turaev Lie bialgebra $(|\hat{T}(H)_{\omega}/\mathbb{K}1|,[\cdot,\cdot]_{\mathrm{gr}},\delta_{\mathrm{gr}})$ is given by $\mathrm{KRV}_{(g,0)}=\operatorname{exp}(\mathfrak{krv}_{(g,0)})$, where
$\mathfrak{krv}_{(g,0)}:=\{g\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})\colon\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)\in\operatorname{Ker}(|\bar{\Delta}_{\omega}|)\}\,.$
Here, $\bar{\Delta}_{\omega}\colon\hat{T}(H)_{\omega}\to\hat{T}(H)_{\omega}\hat{\otimes}\hat{T}(H)_{\omega}$ is the reduced coproduct: $\bar{\Delta}_{\omega}(x)=\Delta_{\omega}(x)-x\otimes 1-1\otimes x$.
We need some lemmas for the proof. The symbol [,[$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the usual commutator, while $[\cdot,\cdot]_{\mathrm{gr}}$ denotes the associated graded of the Goldman bracket.
Lemma 6.3.
Fix a symplectic basis $(x_{i},y_{i})_{1\leq i\leq g}$ of $H$ as before. Let $z\in T(H)_{\omega}$ be a primitive non-zero monomial in $(x_{i},y_{i})_{1\leq i\leq g}$ with $\deg(z)\leq 2$ (i.e., not of the form $(z^{\prime})^{m}$ for $m\geq 2$), and $b\in T(H)_{\omega}$ a homogeneous element. If $|bz^{\ell}|=0$ for a sufficiently large $\ell$, we have $b\in[z,T(H)_{\omega}]$ and therefore $|b|=0$.
Proof. This is a version of Lemma A.3 in [AKKN23], so we basically follow them. If $g=1$, the spaces $T(H)_{\omega}$ and $|T(H)_{\omega}|$ are both isomorphic to the (commutative) polynomial algebra $S(H)$ so we have $b=0$. Now assume $g\geq 2$. We prove by induction on $\deg(b)$. If $\deg(b)=0$, we have $b=0$ since $b$ is just a scalar. Suppose $\deg(b)\geq 1$.
Case of $\deg(z)=1$. We can assume $z=x_{1}$ by applying a graded algebra automorphism of $T(H)_{\omega}$ that sends $z$ to $x_{1}$. Let $\tilde{b}\in T(H)$ be a lift of $b$ such that there is no sequence $y_{g}x_{g}$ appearing in $\tilde{b}$. Then, we have $\operatorname{irr}(|\tilde{b}x_{1}^{\ell}|)=0$ by the assumption $g\geq 2$. Furthermore, each term in $|\tilde{b}x_{1}^{\ell}|$ is not of the form $|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|$ for any $s,t\geq 1$ since $\ell$ is sufficiently large. Then, we have $|\tilde{b}x_{1}^{\ell}|\in\mathrm{Span}_{\mathbb{K}}(X)$. By Theorem 5.22, we obtain $|\tilde{b}x_{1}^{\ell}|=0$ in $|T(H)|$.
Case of $\deg(z)=2$. We can assume $z=x_{1}x_{2}$ or $x_{1}y_{1}$ for the same reason as above. Let $\tilde{b}\in T(H)$ be a lift of $b$ such that there is no sequence $y_{g}x_{g}$ appearing in $\tilde{b}$. Then, we have $\operatorname{irr}(|\tilde{b}z^{\ell}|)=0$ by the assumption $g\geq 2$. Furthermore, each term in $|\tilde{b}z^{\ell}|$ is not of the form $|x_{g}^{s-1}y_{g}^{t-1}y_{g-1}x_{g-1}|$ for any $s,t\geq 1$ since $\ell$ is sufficiently large. Then, we have $|\tilde{b}z^{\ell}|\in\mathrm{Span}_{\mathbb{K}}(X)$. By Theorem 5.22, we obtain $|\tilde{b}z^{\ell}|=0$ in $|T(H)|$.
Therefore, in any case, we have $|\tilde{b}z^{\ell}|=0$ in $|T(H)|$. Now uniquely express $\tilde{b}$ as $\tilde{b}=z\tilde{b}_{1}+\tilde{b}_{2}z+\tilde{b}_{3}$ where $\tilde{b}_{i}\in T(H)$, the terms in $\tilde{b}_{2}$ cannot divided by $z$ from left, and the terms in $\tilde{b}_{3}$ cannot be divided by $z$ from neither side. We have
$0=|\tilde{b}_{1}z^{\ell+1}+\tilde{b}_{2}z^{\ell+1}+\tilde{b}_{3}z^{\ell}|.$
The first and second terms have $(\ell+1)$ consecutive $z$’s in their terms, while none of $|\tilde{b}_{3}z^{\ell}|$ does by the assumption. Therefore we have $|\tilde{b}_{3}z^{\ell}|=0$ and hence $|(\tilde{b}_{1}+\tilde{b}_{2})z^{\ell+1}|=0$. Since $\ell$ is sufficiently large and $\tilde{b}_{3}$ cannot be divided by $z$ from neither side, we have $\tilde{b}_{3}=0$. In addition, by the induction hypothesis, we have $\tilde{b}_{1}+\tilde{b}_{2}=[z,\tilde{b}_{4}]$ for some $\tilde{b}_{4}\in T(H)$. We obtain
$\tilde{b}=z\tilde{b}_{1}+([z,\tilde{b}_{4}]-\tilde{b}_{1})z=[z,\tilde{b}_{1}+\tilde{b}_{4}z].$
Therefore, we have $b=[z,b_{1}+b_{4}z]$; this completes the proof.∎
Lemma 6.4.
Putting $\mathfrak{g}=|\hat{T}(H)_{\omega}/\mathbb{K}1|$ regarded as a Lie algebra, we have $(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}=0$.
Proof. Since the defining equations of the space $(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}$ are linear, we can assume $\mathbb{K}$ to be algebraically closed. Furthermore, since the space $(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}$ is graded, we can work in the non-completed $T(H)_{\omega}$ without any problem. Then, it is reduced to Crawley–Boevey’s result that the quiver variety is generically symplectic: let $Q$ be the quiver with one vertex and $g$ arrows. Then, the preprojective algebra $\Pi$ associated with $Q$ is isomorphic to $T(H)_{\omega}$ as a $\mathbb{K}$-algebra. Since we know the Lie algebra structure on $|T(H)_{\omega}|$ coincides with the necklace Lie bracket on $|\Pi|$ (see Section 6 of [KK16], for example), Theorem 8.6.1 (ii) of [CBEG07] implies $Z(|T(H)_{\omega}|)=Z(|\Pi|)=|\mathbb{K}1|$.
Now take a homogeneous element $x$ of $(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}$ with $\deg(x)=r$. Denote by $\tilde{x}\in|T(H)_{\omega}|^{(\geq 1)\,\otimes 2}$ the unique lift of $x$ along the natural projection $|\hat{T}(H)_{\omega}|^{\otimes 2}\twoheadrightarrow\mathfrak{g}^{\otimes 2}$. The condition $x\in(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}$ is equivalent to $[\tilde{y},\tilde{x}]_{\mathrm{gr}}\in|1|\otimes|T(H)_{\omega}|+|T(H)_{\omega}|\otimes|1|$ for all homogeneous $y\in\mathfrak{g}$ and its unique lift $\tilde{y}\in|T(H)_{\omega}|^{(\geq 1)}$. Put $\tilde{x}=\sum_{0<i<r}\sum_{\lambda\in\Lambda_{i}}\tilde{x}_{i,\lambda}^{\prime}\otimes w_{i,\lambda}$ where $\tilde{x}_{i,\lambda}^{\prime}\in|T(H)_{\omega}|$ with $\deg(x_{i,\lambda}^{\prime})=i$ and $w_{i,\lambda}$ is an element of some fixed homogeneous basis of $|T(H)_{\omega}|$. We have
$[\tilde{y},\tilde{x}]_{\mathrm{gr}}=\sum_{0<i<r}\sum_{\lambda\in\Lambda_{i}}([\tilde{y},\tilde{x}_{i,\lambda}^{\prime}]_{\mathrm{gr}}\otimes w_{i,\lambda}+\tilde{x}_{i,\lambda}^{\prime}\otimes[\tilde{y},w_{i,\lambda}]_{\mathrm{gr}}).$
If $\deg(y)\geq 2$, we have $\deg([\tilde{y},\tilde{x}_{i,\lambda}^{\prime}]_{\mathrm{gr}})=\deg(y)+i-2\geq i\geq 1$ and, similarly, $\deg([\tilde{y},w_{i.\lambda}]_{\mathrm{gr}})\geq 1$ by Proposition 2.8. Therefore, in this case, we obtain $[\tilde{y},\tilde{x}]_{\mathrm{gr}}=0$ for any $y$ with $\deg(y)\geq 2$.
Suppose $\tilde{x}\neq 0$. Take the largest $i$ such that $\tilde{x}_{i,\lambda}^{\prime}\neq 0$ for some $\lambda$. Then, if $\deg(y)\geq 3$, we have $[\tilde{y},\tilde{x}_{i,\lambda}^{\prime}]_{\mathrm{gr}}=0$ for all $\lambda\in\Lambda_{i}$ since the map $[\tilde{y},\cdot\,]_{\mathrm{gr}}$ have degree $\geq 1$ and we took $i$ to be the largest. Let $u_{i,\lambda}^{\prime}\in\operatorname{Der}_{\mathbb{K}}(T(H)_{\omega})$ be a representative of $\sigma_{\mathrm{gr}}(\tilde{x}_{i,\lambda}^{\prime})\in\mathrm{HH}^{1}(\Pi)$. Then, $[\tilde{y},\tilde{x}_{i,\lambda}^{\prime}]_{\mathrm{gr}}=0$ is equivalent to $|u_{i,\lambda}^{\prime}(\tilde{y})|=0$. In particular, for any monomial $a\in\mathfrak{g}$ and $\ell\geq 3$ we have
$\displaystyle 0=|u_{i,\lambda}^{\prime}(\tilde{a}^{\ell})|=\ell|u_{i,\lambda}^{\prime}(\tilde{a})\tilde{a}^{\ell-1}|.$
By Lemma 6.3, for $a\in\mathfrak{g}$ with $\deg(a)\leq 2$, we have $|u_{i,\lambda}^{\prime}(\tilde{a})|=0$. In fact, if $\tilde{a}$ is primitive, we can directly apply Lemma 6.3. If $\tilde{a}=|z^{2}|$ for some $z\in H$, we have $u^{\prime}_{i,\lambda}(z)=[z,b]$ for some $b\in T(H)_{\omega}$ again by Lemma 6.3 and hence
$|u^{\prime}_{i,\lambda}(z^{2})|=2|zu^{\prime}_{i,\lambda}(z)|=2|z[z,b]|=0.$
In conclusion, we have $|u_{i,\lambda}^{\prime}(\tilde{y})|=0$ for any $y\in\mathfrak{g}$, which, in turn, implies $\tilde{x}_{i,\lambda}^{\prime}$ is in the centre $Z(|T(H)_{\omega}|)$. This is a contradiction since $\deg(x^{\prime}_{i,\lambda})\geq 1$ and $\tilde{x}^{\prime}_{i,\lambda}\neq 0$ for some $\lambda$. Therefore, we have $\tilde{x}=0$ and hence $x=0$. ∎
Lemma 6.5.
Any $F\in\operatorname{Aut}^{+}_{\mathrm{Hopf}}(\hat{T}(H)_{\omega})$ preserves the graded Hamiltonian flow $\sigma_{\mathrm{gr}}\colon|\hat{T}(H)_{\omega}|\to\mathrm{HH}^{1}(\hat{T}(H)_{\omega})$.
Proof. We have $F=\exp(f)$ for some $f\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$. We show that any element $\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$ can be lifted to $\operatorname{Der}_{\omega}^{+}(\hat{L}(H))=\{\tilde{f}\in\operatorname{Der}^{+}(\hat{L}(H)):\tilde{f}(\omega)=0\}$. Assume $f$ is homogeneous and take any lift $\tilde{f}$ of $f$ to $\operatorname{Der}^{+}(\hat{L}(H))$. We are done if $\tilde{f}(\omega)=0$, but it is not the case in general; instead, $\tilde{f}(\omega)$ is an element of the ideal $[H,[H,\cdots[H,\omega]\cdots]]$ generated by $\omega$. Since $\deg(f(\omega))\geq 1+\deg(\omega)=3$, we can write $\tilde{f}(\omega)=\sum_{i}[x_{i},a_{i}]+[y_{i},b_{i}]$ with $a_{i},b_{i}\in[H,[H,\cdots[H,\omega]\cdots]]$. Now define $f^{\prime}\in\operatorname{Der}^{+}(\hat{L}(H))$ by $f^{\prime}(x_{i})=\tilde{f}(x_{i})+b_{i}$ and $f^{\prime}(y_{i})=\tilde{f}(y_{i})-a_{i}$. Then, $f^{\prime}$ is also a lift of $f$, and
$\displaystyle f^{\prime}(\omega)$ $\displaystyle=\sum_{i}[f^{\prime}(x_{i}),y_{i}]+[x_{i},f^{\prime}(y_{i})]$
$\displaystyle=\tilde{f}(\omega)+\sum_{i}[b_{i},y_{i}]+[x_{i},-a_{i}]$
$\displaystyle=\sum_{i}[x_{i},a_{i}]+[y_{i},b_{i}]+[b_{i},y_{i}]+[x_{i},-a_{i}]$
$\displaystyle=0.$
Therefore, $F\in\operatorname{Aut}^{+}_{\mathrm{Hopf}}(\hat{T}(H)_{\omega})$ can also be lifted to an element of $\operatorname{Aut}^{+}_{\mathrm{Hopf},\omega}(\hat{T}(H))$. Now that Remark 6.10 of [AKKN23] says $\operatorname{Aut}^{+}_{\mathrm{Hopf},\omega}(\hat{T}(H))$ preserves $\sigma_{\mathrm{gr}}$, the claim immediately follows.∎
Remark 6.6.
The above lemma also follows from the fact that $\sigma_{\mathrm{gr}}$ is a composition
$|\hat{T}(H)_{\omega}|=\mathrm{HH}_{0}(\hat{T}(H)_{\omega})\xrightarrow{\mathsf{B}}\mathrm{HH}_{1}(\hat{T}(H)_{\omega})\cong\mathrm{HH}^{1}(\hat{T}(H)_{\omega})$
where $\mathsf{B}$ is Connes’ differential, and the isomorphism is given by the Poincaré duality. Since $\mathsf{B}$ is functorial, it commutes with $F$. The Poincaré duality is given by the cap product with the fundamental class in $\mathrm{HH}_{2}(\hat{T}(H)_{\omega})$, which is preserved by $F$ since we have assumed $\mathrm{gr}(F)=\operatorname{id}$.
The connection $\nabla^{\prime}\!_{\bullet,H}$ induces a flat homological connection $\nabla\!_{\bullet,H}$ on $0\to Q_{1}\to Q_{0}\to\Omega^{1}\hat{T}(H)_{\omega}\to 0$, where the resolution is given by
$Q_{0}=\hat{T}(H)_{\omega}\otimes H\otimes\hat{T}(H)_{\omega}\;\textrm{ and }\;Q_{1}=\hat{T}(H)_{\omega}\otimes\mathbb{K}\omega\otimes\hat{T}(H)_{\omega}.$
Lemma 6.7.
The associated graded of the Turaev cobracket is equal to the composition $\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}$.
Proof. This is obtained by taking the associated graded of Theorem 8.2 of [Tan24b]. ∎
Proof of Theorem 6.2. We basically imitate the proof in [AKKN23], which is for the surface with non-empty boundary.
Suppose that $G\in\operatorname{Aut}^{+}(\hat{L}(H)_{\omega})$ induces an automorphism of the graded Goldman–Turaev Lie bialgebra. Since $G$ automatically preserves the Lie bracket by Lemma 6.5, this is equivalent to that $G$ only preserves the Lie cobracket. By Lemma 6.7, this is equivalent to
$G\circ\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}-\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}\circ G=0.$
Since $G\in\operatorname{Aut}^{+}(\hat{L}(H)_{\omega})$, we can put $g=\log(G)\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$. Then, the above is further equivalent to
$g\circ\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}-\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}\circ g=0.$
By Lemma 6.5, we have $\operatorname{Ad}_{G}\circ\,\sigma_{\mathrm{gr}}=\sigma_{\mathrm{gr}}\circ G$ on $\mathfrak{g}$, which in turn implies $\operatorname{ad}_{g}\circ\,\sigma_{\mathrm{gr}}=\sigma_{\mathrm{gr}}\circ g$. Therefore, we have
$g\circ\operatorname{Div}^{\nabla_{\bullet,H}}\circ\,\sigma_{\mathrm{gr}}-\operatorname{Div}^{\nabla_{\bullet,H}}\circ\operatorname{ad}_{g}\circ\,\sigma_{\mathrm{gr}}=0.$
By the definition of the infinitesimal adjoint action, we have $g\circ\operatorname{Div}^{\nabla_{\bullet,H}}-\operatorname{Div}^{\nabla_{\bullet,H}}\circ\operatorname{ad}_{g}=\operatorname{Tr}(\operatorname{ad}_{g}\!\nabla_{\bullet,H})$, which is equal to $d(\operatorname{Div}^{\nabla_{\bullet,H}}(g))$ by Corollary A.6 from [Tan25]. Hence, for any $a\in\mathfrak{g}$ and $u=\sigma_{\mathrm{gr}}(a)$, we have
$0=u(\operatorname{Div}^{\nabla_{\bullet,H}}(g))=[a,\operatorname{Div}^{\nabla_{\bullet,H}}(g)]_{\mathrm{gr}}.$
Therefore, $G$ being a formality isomorphism is equivalent to $\operatorname{Div}^{\nabla_{\bullet,H}}(g)\in(\mathfrak{g}\otimes\mathfrak{g})^{\mathfrak{g}}$. By Lemma 6.4, this is equivalent to $\operatorname{Div}^{\nabla_{\bullet,H}}(g)=0$ in $\mathfrak{g}^{\otimes 2}$.
Now we have $\operatorname{Div}^{\nabla_{\bullet,H}}(g)=|\tilde{\Delta}_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))$ by Section 6 of [Tan24b]. Since the antipode is bijective, it is equivalent to $|\Delta_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))=0\in\mathfrak{g}^{\otimes 2}$. In addition, $\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)$ has no constant term since the connection and hence the divergence both have degree 0 while $g$ has a positive degree. Therefore, using a canonical map $\mathfrak{g}\xrightarrow{\cong}|\hat{T}(H)_{\omega}|^{(\geq 1)}\subset|\hat{T}(H)_{\omega}|$, the equality is restated as
$|\Delta_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))-|1|\otimes\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)-\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)\otimes|1|=0\;\textrm{ in }|\hat{T}(H)_{\omega}|^{\otimes 2}.$
This is exactly the statement of the theorem.∎
7. The Kernel of the Reduced Coproduct
In this section, we determine the space $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)$ appeared in Theorem 6.2 in low degrees. Let $L(H)$ be the free Lie algebra over $H$ (without completion) and $L(H)_{\omega}$ the quotient of $L(H)$ by the ideal generated by $\omega$. We have $T(H)_{\omega}=UL(H)_{\omega}$.
Notations. The superscript (d) denotes the degree $d$ part for $d\in\mathbb{Z}$. For $\mathbb{K}$-subspaces $Z_{1},Z_{2}$ of a $\mathbb{K}$-algebra $A$, we denote by $|Z_{1}Z_{2}|$ the image of the composite
$Z_{1}\otimes Z_{2}\xhookrightarrow{\mathrm{incl}^{\otimes 2}}A\otimes A\xrightarrow{\mathrm{mult}}A\twoheadrightarrow|A|.$
The main result of this section is the following.
Proposition 7.1.
In low degrees, $\operatorname{Ker}(|\bar{\Delta}_{\omega}|\colon|T(H)_{\omega}|\to|T(H)_{\omega}|^{\otimes 2})$ is given explicitly by
$\displaystyle\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(1)}$ $\displaystyle=H,$
$\displaystyle\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(2)}$ $\displaystyle=0,$
$\displaystyle\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(3)}$ $\displaystyle=|HL(H)_{\omega}^{(2)}|\cong\wedge^{3}H/|H\omega|,$
for any $g\geq 1$, and $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(4)}=|HL(H)_{\omega}^{(3)}|$ for $g\neq 2$.
Remark 7.2.
For $g=2$, we have a strict inclusion $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(4)}\supset|HL(H)_{\omega}^{(3)}|$.
7.1 The Case of a Free Associative Algebra
We will start with the case of the free associative algebra $T(H)$ and later apply the result to the case of the quotient algebra $T(H)_{\omega}$.
Lemma 7.3.
In low degrees, $\operatorname{Ker}(|\bar{\Delta}|\colon|T(H)|\to|T(H)|^{\otimes 2})$ is given explicitly by
$\displaystyle\operatorname{Ker}(|\bar{\Delta}|)^{(1)}$ $\displaystyle=H,$
$\displaystyle\operatorname{Ker}(|\bar{\Delta}|)^{(2)}$ $\displaystyle=0,$
$\displaystyle\operatorname{Ker}(|\bar{\Delta}|)^{(3)}$ $\displaystyle=|HL(H)^{(2)}|\cong\wedge^{3}H,\textrm{ and}$
$\displaystyle\operatorname{Ker}(|\bar{\Delta}|)^{(4)}$ $\displaystyle=|HL(H)^{(3)}|$
for any $g\geq 1$.
Proof. The degree 1 part is easily checked. In degree 2, the reduced coproduct reads
$\bar{\Delta}\colon S^{2}(H)=|HH|\to H\otimes H\colon|xy|\mapsto|x|\otimes|y|+|y|\otimes|x|.$
This is an isomorphism onto the image.
Next, the inclusion $\operatorname{Ker}(|\bar{\Delta}|)^{(d)}\supset|HL(H)^{(d-1)}|$ for $d\geq 3$ is checked by the following calculation: for $x\in H$ and $\ell\in L$,
$|\Delta(x\ell)|=|\Delta(x)\Delta(\ell)|=|(x\otimes 1+1\otimes x)(\ell\otimes 1+1\otimes\ell)|=|x\ell\otimes 1+x\otimes\ell+\ell\otimes x+1\otimes x\ell|.$
Since $|L(H)^{(d-1)}|=0$ for $d\geq 3$, $|x\otimes\ell+\ell\otimes x|=|x|\otimes|\ell|+|\ell|\otimes|x|$ vanishes if $\deg(\ell)\geq 2$.
Considering the multi-grading, we show the converse is true for $d=3,4$. Take a $\mathbb{K}$-basis $(v_{i})_{1\leq i\leq 2g}$ of $H$. The multi-degree is given by the decomposition $T(H)=\bigoplus_{\mathbf{d}\in\mathbb{Z}^{2g}}T(H)_{\mathbf{d}}$ specified by $\deg(v_{i})=e_{i}$. Here $e_{i}$ is the $i$-th standard unit vector of $\mathbb{Z}^{2g}$. The homogeneous part $T(H)_{\mathbf{d}}$ is zero if any coefficient of $e_{i}$ is negative. The coproduct $\Delta$ and hence the reduced $\bar{\Delta}$ are degree 00 0 with respect to this grading, so we only have to compute the kernel for each fixed multi-degree. The total degree map is defined by $|\cdot|\colon\mathbb{Z}^{2g}\to\mathbb{Z}\colon e_{i}\mapsto 1$.
In degree 3, only possible multi-degrees are, up to permutations of $e_{i}$’s,
  1. (i)
    $(3,0,\dotsc,0)$;
  2. (ii)
    $(2,1,0,\dotsc,0)$; or
  3. (iii)
    $(1,1,1,0,\dotsc,0)$.
We compute the kernel in each case and show that it is contained in $|HL(H)^{(2)}|$.
  1. (i)
    The only mononial with multi-degree $(3,0,\dotsc,0)$ is $|v_{1}^{3}|$. Then, we have $0=\bar{\Delta}(c|v_{1}^{3}|)=3c|v_{1}|\otimes|v_{1}^{2}|$ for $c\in\mathbb{K}$, which implies $c=0$.
  2. (ii)
    Similarly, for $c\in\mathbb{K}$, we have $0=\bar{\Delta}(c|v_{1}^{2}v_{2}|)=2c|v_{1}|\otimes|v_{1}v_{2}|+c|v_{2}|\otimes|v_{1}^{2}|$. Again $c=0$.
  3. (iii)
    For $c,c^{\prime}\in\mathbb{K}$, we have
    00$\displaystyle$ 0 $\displaystyle=\bar{\Delta}(c|v_{1}v_{2}v_{3}|+c^{\prime}|v_{1}v_{3}v_{2}|)$
    $\displaystyle=c(|v_{1}|\otimes|v_{2}v_{3}|+|v_{2}|\otimes|v_{1}v_{3}|+|v_{3}|\otimes|v_{1}v_{2}|)+c^{\prime}(|v_{1}|\otimes|v_{3}v_{2}|+|v_{3}|\otimes|v_{1}v_{2}|+|v_{2}|\otimes|v_{1}v_{3}|).$
    Therefore, we obtain $c+c^{\prime}=0$, and we have $c|v_{1}v_{2}v_{3}|+c^{\prime}|v_{1}v_{3}v_{2}|=c|v_{1}[v_{2},v_{3}]|\in|HL(H)^{(2)}|$.
This completes the case of degree 3.
Now, we deal with the degree 44 44 case. Only possible multi-degrees with $|\mathbf{d}|=4$ are, up to permutations of $e_{i}$’s, one of the following:
  1. (i)
    $(4,0,\dotsc,0)$;
  2. (ii)
    $(3,1,0,\dotsc,0)$;
  3. (iii)
    $(2,2,0,\dotsc,0)$;
  4. (iv)
    $(2,1,1,0,\dotsc,0)$; or
  5. (v)
    $(1,1,1,1,0,\dotsc,0)$.
We compute the kernel in each case and show that it is contained in $|HL(H)^{(3)}|$.
  1. (i)
    The only mononial with multi-degree $(4,0,\dotsc,0)$ is $v_{1}^{4}$. We put $a=cv_{1}^{4}$ with $c\in\mathbb{K}$. Then, we have
    $0=|4c\,v_{1}\otimes v_{1}^{3}+6c\,v_{1}^{2}\otimes v_{1}^{2}|.$
    by seeing the degree $(1,3)$ part of $|\bar{\Delta}(a)|$. Therefore, $|a|\in\operatorname{Ker}(|\bar{\Delta}|)$ implies $c=0$.
  2. (ii)
    In multi-degree $(3,1,0,\dotsc,0)$, it is enough to consider $a=cv_{1}^{3}v_{2}$. Then,
    $0=|3cv_{1}\otimes v_{1}^{2}v_{2}+cv_{2}\otimes v_{1}^{3}|.$
    Therefore, $|a|\in\operatorname{Ker}(|\bar{\Delta}|)$ implies $c=0$.
  3. (iii)
    In multi-degree $(2,2,0,\dotsc,0)$, it is enough to consider the case $a=c_{1}v_{1}^{2}v_{2}^{2}+c_{2}v_{1}v_{2}v_{1}v_{2}$. Then,
    $0=|2c_{1}(v_{1}\otimes v_{1}v_{2}^{2}+v_{2}\otimes v_{1}^{2}v_{2})+2c_{2}(v_{1}\otimes v_{1}v_{2}^{2}+v_{2}\otimes v_{1}^{2}v_{2})|.$
    Therefore, $|a|\in\operatorname{Ker}(|\bar{\Delta}|)$ implies $c_{1}+c_{2}=0$. We have
    $\displaystyle|a|$ $\displaystyle=c_{1}|v_{1}(v_{1}v_{2}^{2}-v_{2}v_{1}v_{2})|$
    $\displaystyle=c_{1}|v_{1}[v_{1},v_{2}]v_{2}|$
    $\displaystyle=\frac{c_{1}}{2}|v_{1}[[v_{1},v_{2}],v_{2}]|.$
    Thus we have $|a|\in|HL(H)^{(3)}|$.
  4. (iv)
    In multi-degree $(2,1,1,0,\dotsc,0)$, it is enough to consider the case $a=c_{1}v_{1}^{2}v_{2}v_{3}+c_{2}v_{1}v_{2}v_{1}v_{3}+c_{3}v_{2}v_{1}^{2}v_{3}$. Then,
    00$\displaystyle$ 0 $\displaystyle=|c_{1}(2v_{1}\otimes v_{1}v_{2}v_{3}+v_{2}\otimes v_{1}^{2}v_{3}+v_{3}\otimes v_{1}^{2}v_{2})$
    $\displaystyle\qquad+c_{2}(v_{1}\otimes v_{2}v_{1}v_{3}+v_{2}\otimes v_{1}^{2}v_{3}+v_{1}\otimes v_{1}v_{2}v_{3}+v_{3}\otimes v_{1}^{2}v_{2})$
    $\displaystyle\qquad+c_{3}(v_{2}\otimes v_{1}^{2}v_{3}+2v_{1}\otimes v_{2}v_{1}v_{3}+v_{3}\otimes v_{1}^{2}v_{2})|.$
    Therefore, $|a|\in\operatorname{Ker}(|\bar{\Delta}|)$ implies $0\displaystyle 2c_{1}+c_{2}=$ 0, $c_{2}+2c_{3}=0$ and $c_{1}+c_{2}+c_{3}=0$. We have
    $|a|=c_{1}|v_{1}^{2}v_{2}v_{3}-2v_{1}v_{2}v_{1}v_{3}+v_{2}v_{1}^{2}v_{3}|=c_{1}|[v_{1},[v_{1},v_{2}]]v_{3}|.$
    Thus we have $|a|\in|HL(H)^{(3)}|$.
  5. (v)
    In multi-degree $(1,1,1,1,0,\dotsc,0)$, it is enough to consider the case $a=c_{1}v_{1}v_{2}v_{3}v_{4}+c_{2}v_{1}v_{3}v_{2}v_{4}+c_{3}v_{2}v_{1}v_{3}v_{4}+c_{4}v_{2}v_{3}v_{1}v_{4}+c_{5}v_{3}v_{1}v_{2}v_{4}+c_{6}v_{3}v_{2}v_{1}v_{4}$. Then,
    00$\displaystyle$ 0 $\displaystyle=|c_{1}(v_{1}\otimes v_{2}v_{3}v_{4}+v_{2}\otimes v_{1}v_{3}v_{4}+v_{3}\otimes v_{1}v_{2}v_{4}+v_{4}\otimes v_{1}v_{2}v_{3})$
    $\displaystyle\qquad+c_{2}(v_{1}\otimes v_{3}v_{2}v_{4}+v_{2}\otimes v_{1}v_{3}v_{4}+v_{3}\otimes v_{1}v_{2}v_{4}+v_{4}\otimes v_{1}v_{3}v_{2})$
    $\displaystyle\qquad+c_{3}(v_{1}\otimes v_{2}v_{3}v_{4}+v_{2}\otimes v_{1}v_{3}v_{4}+v_{3}\otimes v_{2}v_{1}v_{4}+v_{4}\otimes v_{2}v_{1}v_{3})$
    $\displaystyle\qquad+c_{4}(v_{1}\otimes v_{2}v_{3}v_{4}+v_{2}\otimes v_{3}v_{1}v_{4}+v_{3}\otimes v_{2}v_{1}v_{4}+v_{4}\otimes v_{2}v_{3}v_{1})$
    $\displaystyle\qquad+c_{5}(v_{1}\otimes v_{3}v_{2}v_{4}+v_{2}\otimes v_{3}v_{1}v_{4}+v_{3}\otimes v_{1}v_{2}v_{4}+v_{4}\otimes v_{3}v_{1}v_{2})$
    $\displaystyle\qquad+c_{6}(v_{1}\otimes v_{3}v_{2}v_{4}+v_{2}\otimes v_{3}v_{1}v_{4}+v_{3}\otimes v_{2}v_{1}v_{4}+v_{4}\otimes v_{3}v_{2}v_{1})|.$
    Therefore we have
    $\displaystyle c_{1}+c_{3}+c_{4}=0,\quad c_{2}+c_{5}+c_{6}=0,$
    $\displaystyle c_{1}+c_{2}+c_{3}=0,\quad c_{4}+c_{5}+c_{6}=0,$
    $\displaystyle c_{1}+c_{2}+c_{5}=0,\quad c_{3}+c_{4}+c_{6}=0,$
    $\displaystyle c_{1}+c_{4}+c_{5}=0,\textrm{ and }c_{2}+c_{3}+c_{6}=0.$
    This boils down to $c_{1}=c_{6}$, $c_{2}=c_{4}$, $c_{3}=c_{5}$ and $c_{1}+c_{2}+c_{3}=0$. Finally, we have
    $\displaystyle|a|$ $\displaystyle=|c_{1}v_{1}v_{2}v_{3}v_{4}+c_{2}v_{1}v_{3}v_{2}v_{4}-(c_{1}+c_{2})v_{2}v_{1}v_{3}v_{4}+c_{2}v_{2}v_{3}v_{1}v_{4}-(c_{1}+c_{2})v_{3}v_{1}v_{2}v_{4}+c_{1}v_{3}v_{2}v_{1}v_{4}|$
    $\displaystyle=|c_{1}[[v_{1},v_{2}],v_{3}]v_{4}+c_{2}[[v_{1},v_{3}],v_{2}]v_{4}|$
    and we obtain $|a|\in|HL(H)^{(3)}|$.
This completes the proof.∎
Remark 7.4.
Lemma 7.3 holds for an arbitrary $\mathbb{K}$-vector space $H$; the proof is identical.
7.2 Proof of Proposition 7.1.
The kernel of the projection $|T(H)|^{\otimes 2}\to|T(H)_{\omega}|^{\otimes 2}$ is equal to $|\langle\omega\rangle|\otimes|T(H)|\oplus|T(H)|\otimes|\langle\omega\rangle|$. Since $|\langle\omega\rangle|^{(\leq 2)}=0$, the intersection with $\operatorname{Im}|\bar{\Delta}|\subset|T(H)|\otimes H+H\otimes|T(H)|$ is zero in degree $d\leq 3$. Hence, the claim follows.
Now, we consider the degree 4 case. We modify the above multi-grading as follows. Take a symplectic basis $(x_{i},y_{i})_{1\leq i\leq g}$ of $H$ as before, and set $v_{2i-1}=x_{i}$ and $v_{2i}=y_{i}$ for $1\leq i\leq g$. The algebra $T(H)_{\omega}=T(H)/\langle\omega\rangle$ is graded by the abelian group $D=\mathbb{Z}^{2g}/\langle e_{1}+e_{2}=e_{3}+e_{4}=\cdots=e_{2g-1}+e_{2g}\rangle$ where $(e_{i})_{1\leq i\leq 2g}$ is the standard basis of $\mathbb{Z}^{2g}$ and $\deg(v_{i})=e_{i}$. Indeed, $\omega$ is homogeneous with respect to the $D$-grading. We have a decomposition $T(H)_{\omega}=\bigoplus_{\mathbf{d}\in D}(T(H)_{\omega})_{\mathbf{d}}$. Denoting the natural map by $p\colon(\mathbb{Z}_{\geq 0})^{2g}\to D$, the homogeneous part $(T(H)_{\omega})_{\mathbf{d}}$ is non-zero if and only if $\mathbf{d}\in D^{+}:=p((\mathbb{Z}_{\geq 0})^{2g})$. The total degree map $|\cdot|\colon\mathbb{Z}^{2g}\to\mathbb{Z}$ descends to $|\cdot|\colon D\to\mathbb{Z}$.
Define the redundancy $R\colon D^{+}\to\mathbb{Z}$ by $R(\mathbf{d})=\sum_{1\leq i\leq g}\mathrm{min}(\lambda_{2i-1},\lambda_{2i})$ with $\mathbf{d}=p\Big{(}\sum_{i}\lambda_{i}e_{i}\Big{)}$. This is well-defined, and we have $2\displaystyle 0\leq R(\mathbf{d})\leq|\mathbf{d}|/$ 2 for any $\mathbf{d}$. Now that we set $|\mathbf{d}|=4$, we only have to inspect the following cases:
  1. (0)
    If $R(\mathbf{d})=0$, we have #$\#p^{-1}(\mathbf{d})=1$. Therefore, the natural map $T(H)_{p^{-1}(\mathbf{d})}\twoheadrightarrow(T(H)_{\omega})_{\mathbf{d}}$ is an isomorphism. The computation of $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)$ is reduced to the case of $T(H)$.
  2. (1)
    If $R(\mathbf{d})=1$, we have #$\#p^{-1}(\mathbf{d})=g$. The inverse image $p^{-1}(\mathbf{d})$ is, up to permutations of pairs $(x_{i},y_{i})\leftrightarrow(x_{j},y_{j})$ and $(x_{i},y_{i})\leftrightarrow(y_{i},-x_{i})$, one of the followings:
    1. (i)
      $\{e_{2i-1}+e_{2i}+2e_{2g-1}:1\leq i\leq g\}$; or
    2. (ii)
      $\{e_{2i-1}+e_{2i}+e_{2g-3}+e_{2g-1}:1\leq i\leq g\}$.
  3. (2)
    If $R(\mathbf{d})=2$, we have #$\#p^{-1}(\mathbf{d})=\frac{g(g+1)}{2}$. There is only one such $\mathbf{d}\in D^{+}$, and the inverse image is given by $p^{-1}(\mathbf{d})=\{e_{2i-1}+e_{2i}+e_{2j-1}+e_{2j}:1\leq i\leq j\leq g\}$.
Since the case $(0)$ is already done, we will deal with the cases $(1)$ and $(2)$ from now on. The case $g=1$ is easy, so we assume $g\geq 3$.
(1-i) It suffices to consider $a=\sum_{1\leq i<g}(c^{i}_{1}x_{i}y_{i}x_{g}^{2}+c^{i}_{2}x_{i}x_{g}y_{i}x_{g}+c^{i}_{3}y_{i}x_{i}x_{g}^{2})+c_{1}^{g}x_{g}^{3}y_{g}$. We may take $c^{1}_{1}=-c_{2}^{1}/2$ by subtracting some scalar multiple of $\omega x_{g}^{2}$ from $a$. Then,
00$\displaystyle$ 0 $\displaystyle=\sum_{1\leq i<g}|c_{1}^{i}(x_{i}\otimes y_{i}x_{g}^{2}+y_{i}\otimes x_{i}x_{g}^{2}+2x_{g}\otimes x_{i}y_{i}x_{g})$
$\displaystyle\qquad\qquad+c_{2}^{i}(x_{i}\otimes x_{g}y_{i}x_{g}+y_{i}\otimes x_{i}x_{g}^{2}+x_{g}\otimes x_{i}y_{i}x_{g}+x_{g}\otimes x_{i}x_{g}y_{i})$
$\displaystyle\qquad\qquad+c_{3}^{i}(x_{i}\otimes y_{i}x_{g}^{2}+y_{i}\otimes x_{i}x_{g}^{2}+2x_{g}\otimes y_{i}x_{i}x_{g})|$
$\displaystyle\qquad+c_{1}^{g}|3x_{g}\otimes x_{g}^{2}y_{g}+y_{g}\otimes x_{g}^{3}|.$
We have to use two rewriting rules, but the only applicable term is $|y_{g-1}x_{g-1}x_{g}|$ in the seventh and tenth terms. By applying $\rho_{2}$ with $s=2$ and $t=1$, we have
$|x_{g}y_{g-1}x_{g-1}|\mapsto|x_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime})|.$
At this point, we can simply compare the coefficients by Theorem 5.22. Therefore, we have
[…]
$\displaystyle c_{1}^{g}=0$
Since we took $c^{1}_{1}=-c_{2}^{1}/2$, we have $0\displaystyle 2c_{1}^{i}+c_{2}^{i}=c_{2}^{i}+2c_{3}^{i}=$ 0 for all $1\leq i<g$. The rest is the same as the case of $T(H)$.
(1-ii) It suffices to consider
$\displaystyle a$ $\displaystyle=\sum_{1\leq i<g-1}(c^{i}_{1}x_{i}y_{i}x_{g-1}x_{g}+c_{2}^{i}x_{i}x_{g-1}y_{i}x_{g}+c_{3}^{i}y_{i}x_{i}x_{g-1}x_{g}+c_{4}^{i}y_{i}x_{g-1}x_{i}x_{g}+c_{5}^{i}x_{g-1}x_{i}y_{i}x_{g}+c_{6}^{i}x_{g-1}y_{i}x_{i}x_{g})$
$\displaystyle\quad+(c_{1}^{g-1}x_{g-1}^{2}y_{g-1}x_{g}+c_{2}^{g-1}x_{g-1}y_{g-1}x_{g-1}x_{g}+c_{3}^{g-1}y_{g-1}x_{g-1}^{2}x_{g})$
$\displaystyle\quad+(c_{1}^{g}x_{g-1}x_{g}^{2}y_{g}+c_{2}^{g}x_{g-1}x_{g}y_{g}x_{g}+c_{3}^{g}x_{g-1}y_{g}x_{g}^{2}).$
We have
00$\displaystyle$ 0 $\displaystyle=\sum_{1\leq i<g-1}|c_{1}^{i}(x_{i}\otimes y_{i}x_{g-1}x_{g}+y_{i}\otimes x_{i}x_{g-1}x_{g}+x_{g-1}\otimes x_{i}y_{i}x_{g}+x_{g}\otimes x_{i}y_{i}x_{g-1})$
$\displaystyle\qquad\qquad\quad+c_{2}^{i}(x_{i}\otimes x_{g-1}y_{i}x_{g}+y_{i}\otimes x_{i}x_{g-1}x_{g}+x_{g-1}\otimes x_{i}y_{i}x_{g}+x_{g}\otimes x_{i}x_{g-1}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{3}^{i}(x_{i}\otimes y_{i}x_{g-1}x_{g}+y_{i}\otimes x_{i}x_{g-1}x_{g}+x_{g-1}\otimes y_{i}x_{i}x_{g}+x_{g}\otimes y_{i}x_{i}x_{g-1})$
$\displaystyle\qquad\qquad\quad+c_{4}^{i}(x_{i}\otimes y_{i}x_{g-1}x_{g}+y_{i}\otimes x_{g-1}x_{i}x_{g}+x_{g-1}\otimes y_{i}x_{i}x_{g}+x_{g}\otimes y_{i}x_{g-1}x_{i})$
$\displaystyle\qquad\qquad\quad+c_{5}^{i}(x_{i}\otimes x_{g-1}y_{i}x_{g}+y_{i}\otimes x_{g-1}x_{i}x_{g}+x_{g-1}\otimes x_{i}y_{i}x_{g}+x_{g}\otimes x_{g-1}x_{i}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{6}^{i}(x_{i}\otimes x_{g-1}y_{i}x_{g}+y_{i}\otimes x_{g-1}x_{i}x_{g}+x_{g-1}\otimes y_{i}x_{i}x_{g}+x_{g}\otimes x_{g-1}y_{i}x_{i})|$
$\displaystyle\qquad+c_{1}^{g-1}|2x_{g-1}\otimes x_{g-1}y_{g-1}x_{g}+y_{g-1}\otimes x_{g-1}^{2}x_{g}+x_{g}\otimes x_{g-1}^{2}y_{g-1}|$
$\displaystyle\qquad+c_{2}^{g-1}|x_{g-1}\otimes y_{g-1}x_{g-1}x_{g}+y_{g-1}\otimes x_{g-1}^{2}x_{g}+x_{g-1}\otimes x_{g-1}y_{g-1}x_{g}+x_{g}\otimes x_{g-1}^{2}y_{g-1}|$
$\displaystyle\qquad+c_{3}^{g-1}|2x_{g-1}\otimes y_{g-1}x_{g-1}x_{g}+y_{g-1}\otimes x_{g-1}^{2}x_{g}+x_{g}\otimes y_{g-1}x_{g-1}^{2}|$
$\displaystyle\qquad+c_{1}^{g}|x_{g-1}\otimes x_{g}^{2}y_{g}+2x_{g}\otimes x_{g-1}x_{g}y_{g}+y_{g}\otimes x_{g-1}x_{g}^{2}|$
$\displaystyle\qquad+c_{2}^{g}|x_{g-1}\otimes x_{g}^{2}y_{g}+x_{g}\otimes x_{g-1}y_{g}x_{g}+y_{g}\otimes x_{g-1}x_{g}^{2}+x_{g}\otimes x_{g-1}x_{g}y_{g}|$
$\displaystyle\qquad+c_{3}^{g}|x_{g-1}\otimes y_{g}x_{g}^{2}+2x_{g}\otimes x_{g-1}y_{g}x_{g}+y_{g}\otimes x_{g-1}x_{g}^{2}|.$
To apply the rewriting rules, we first search for the sequence $y_{g}x_{g}$, which appears in $c_{2}^{g}x_{g}\otimes x_{g-1}y_{g}x_{g}$ and $c_{3}^{g}\cdot 2x_{g}\otimes x_{g-1}y_{g}x_{g}$. The monomial $x_{g-1}y_{g}x_{g}$ is rewritten into $x_{g-1}(x_{g}y_{g}+\omega^{\prime})$. Next, we search for the cyclic monomial $|x_{g}y_{g-1}x_{g-1}|$ or $|y_{g}y_{g-1}x_{g-1}|$. The latter does not appear since it has multi-degree $(0,\dotsc,0,1,1,0,1)$, which is impossible. The former appears in $c_{2}^{g-1}x_{g-1}\otimes y_{g-1}x_{g-1}x_{g}$ and $c_{3}^{g-1}\cdot 2x_{g-1}\otimes y_{g-1}x_{g-1}x_{g}$. After the rewriting, we can compare the coefficients; we have
[…]
$\displaystyle x_{g-1}:(c_{1}^{g-1}+c_{2}^{g-1}+c_{3}^{g-1})=0$ ,
$\displaystyle x_{g-1}:(c_{1}^{g}+c_{2}^{g}+c_{3}^{g})=0$ ,
[…]
From this, we have $c_{2}^{i}=c_{4}^{i}$ for all $1\displaystyle 1\leq i<g-$ 1. Since we can freely take $c_{1}^{1}$ and $c_{1}^{6}$ by subtracting $\omega x_{g-1}x_{g}$ and $x_{g-1}\omega x_{g}$, respectively, we set $c_{1}^{1}+c_{4}^{1}+c_{5}^{1}=c_{2}^{1}+c_{3}^{1}+c_{6}^{1}=0$. Then, we have $c_{2}^{g}+2c_{3}^{g}=2c_{1}^{c}+c_{2}^{g}=0$. Also we have $c_{1}^{1}+c_{2}^{1}+c_{5}^{1}=0$ since $c_{2}^{1}=c_{4}^{1}$, so we get $c_{2}^{g-1}+2c_{3}^{g-1}=0$ and $c_{1}^{i}+c_{2}^{i}+c_{5}^{i}=c_{3}^{i}+c_{4}^{i}+c_{6}^{i}=0$ for all $i$. The rest is the same as the case of $T(H)$.
(2) It suffices to consider
$\displaystyle a$ $\displaystyle=\sum_{1\leq i\leq g}(c_{1}^{i}x_{i}^{2}y_{i}^{2}+c_{2}^{i}x_{i}y_{i}x_{i}y_{i})$
$\displaystyle\qquad+\sum_{1\leq i<j\leq g}(c_{1}^{ij}x_{i}y_{i}x_{j}y_{j}+c_{2}^{ij}x_{i}y_{i}y_{j}x_{j}+c_{3}^{ij}x_{i}x_{j}y_{i}y_{j}+c_{4}^{ij}x_{i}x_{j}y_{j}y_{i}+c_{5}^{ij}x_{i}y_{j}y_{i}x_{j}+c_{6}^{ij}x_{i}y_{j}x_{j}y_{i}).$
We have
00$\displaystyle$ 0 $\displaystyle=\sum_{1\leq i\leq g}|c_{1}^{i}(2x_{i}\otimes x_{i}y_{i}^{2}+2y_{i}\otimes x_{i}^{2}y_{i})+c_{2}^{i}(2x_{i}\otimes x_{i}y_{i}^{2}+2y_{i}\otimes x_{i}^{2}y_{i})|$
$\displaystyle+\sum_{1\leq i<j\leq g}|c_{1}^{ij}(x_{i}\otimes y_{i}x_{j}y_{j}+y_{i}\otimes x_{i}x_{j}y_{j}+x_{j}\otimes x_{i}y_{i}y_{j}+y_{j}\otimes x_{i}y_{i}x_{j})$
$\displaystyle\qquad\qquad\quad+c_{2}^{ij}(x_{i}\otimes y_{i}y_{j}x_{j}+y_{i}\otimes x_{i}y_{j}x_{j}+y_{j}\otimes x_{i}y_{i}x_{j}+x_{j}\otimes x_{i}y_{i}y_{j})$
$\displaystyle\qquad\qquad\quad+c_{3}^{ij}(x_{i}\otimes x_{j}y_{i}y_{j}+x_{j}\otimes x_{i}y_{i}y_{j}+y_{i}\otimes x_{i}x_{j}y_{j}+y_{j}\otimes x_{i}x_{j}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{4}^{ij}(x_{i}\otimes x_{j}y_{j}y_{i}+x_{j}\otimes x_{i}y_{j}y_{i}+y_{j}\otimes x_{i}x_{j}y_{i}+y_{i}\otimes x_{i}x_{j}y_{j})$
$\displaystyle\qquad\qquad\quad+c_{5}^{ij}(x_{i}\otimes y_{j}y_{i}x_{j}+y_{j}\otimes x_{i}y_{i}x_{j}+y_{i}\otimes x_{i}y_{j}x_{j}+x_{j}\otimes x_{i}y_{j}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{6}^{ij}(x_{i}\otimes y_{j}x_{j}y_{i}+y_{j}\otimes x_{i}x_{j}y_{i}+x_{j}\otimes x_{i}y_{j}y_{i}+y_{i}\otimes x_{i}y_{j}x_{j})|$
$\displaystyle\xmapsto{\rho,\rho_{2}}\sum_{1\leq i\leq g}|c_{1}^{i}(2x_{i}\otimes x_{i}y_{i}^{2}+2y_{i}\otimes x_{i}^{2}y_{i})+c_{2}^{i}(2x_{i}\otimes x_{i}y_{i}^{2}+2y_{i}\otimes x_{i}^{2}y_{i})|$
$\displaystyle+\sum_{1\leq i<j<g}|c_{1}^{ij}(x_{i}\otimes y_{i}x_{j}y_{j}+y_{i}\otimes x_{i}x_{j}y_{j}+x_{j}\otimes x_{i}y_{i}y_{j}+y_{j}\otimes x_{i}y_{i}x_{j})$
$\displaystyle\qquad\qquad\quad+c_{2}^{ij}(x_{i}\otimes y_{i}y_{j}x_{j}+y_{i}\otimes x_{i}y_{j}x_{j}+y_{j}\otimes x_{i}y_{i}x_{j}+x_{j}\otimes x_{i}y_{i}y_{j})$
$\displaystyle\qquad\qquad\quad+c_{3}^{ij}(x_{i}\otimes x_{j}y_{i}y_{j}+x_{j}\otimes x_{i}y_{i}y_{j}+y_{i}\otimes x_{i}x_{j}y_{j}+y_{j}\otimes x_{i}x_{j}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{4}^{ij}(x_{i}\otimes x_{j}y_{j}y_{i}+x_{j}\otimes x_{i}y_{j}y_{i}+y_{j}\otimes x_{i}x_{j}y_{i}+y_{i}\otimes x_{i}x_{j}y_{j})$
$\displaystyle\qquad\qquad\quad+c_{5}^{ij}(x_{i}\otimes y_{j}y_{i}x_{j}+y_{j}\otimes x_{i}y_{i}x_{j}+y_{i}\otimes x_{i}y_{j}x_{j}+x_{j}\otimes x_{i}y_{j}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{6}^{ij}(x_{i}\otimes y_{j}x_{j}y_{i}+y_{j}\otimes x_{i}x_{j}y_{i}+x_{j}\otimes x_{i}y_{j}y_{i}+y_{i}\otimes x_{i}y_{j}x_{j})|$
$\displaystyle+\sum_{1\leq i<g-1}|c_{1}^{ig}(x_{i}\otimes y_{i}x_{g}y_{g}+y_{i}\otimes x_{i}x_{g}y_{g}+x_{g}\otimes x_{i}y_{i}y_{g}+y_{g}\otimes x_{i}y_{i}x_{g})$
$\displaystyle\qquad\qquad\quad+c_{2}^{ig}(x_{i}\otimes y_{i}(x_{g}y_{g}+\omega^{\prime})+y_{i}\otimes x_{i}(x_{g}y_{g}+\omega^{\prime})+y_{g}\otimes x_{i}y_{i}x_{g}+x_{g}\otimes x_{i}y_{i}y_{g})$
$\displaystyle\qquad\qquad\quad+c_{3}^{ig}(x_{i}\otimes y_{i}(x_{g}y_{g}+\omega^{\prime})+x_{g}\otimes x_{i}y_{i}y_{g}+y_{i}\otimes x_{i}x_{g}y_{g}+y_{g}\otimes x_{i}x_{g}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{4}^{ig}(x_{i}\otimes x_{g}y_{g}y_{i}+x_{g}\otimes x_{i}y_{g}y_{i}+y_{g}\otimes x_{i}x_{g}y_{i}+y_{i}\otimes x_{i}x_{g}y_{g})$
$\displaystyle\qquad\qquad\quad+c_{5}^{ig}(x_{i}\otimes y_{g}y_{i}x_{g}+y_{g}\otimes x_{i}y_{i}x_{g}+y_{i}\otimes x_{i}(x_{g}y_{g}+\omega^{\prime})+x_{g}\otimes x_{i}y_{g}y_{i})$
$\displaystyle\qquad\qquad\quad+c_{6}^{ig}(x_{i}\otimes(x_{g}y_{g}+\omega^{\prime})y_{i}+y_{g}\otimes x_{i}x_{g}y_{i}+x_{g}\otimes x_{i}y_{g}y_{i}+y_{i}\otimes x_{i}(x_{g}y_{g}+\omega^{\prime}))|$
$\displaystyle+|c_{1}^{g-1,g}(x_{g-1}\otimes y_{g-1}x_{g}y_{g}+y_{g-1}\otimes x_{g-1}x_{g}y_{g}+x_{g}\otimes x_{g-1}y_{g-1}y_{g}+y_{g}\otimes x_{g-1}y_{g-1}x_{g})$
$\displaystyle\quad+c_{2}^{g-1,g}(x_{g-1}\otimes y_{g-1}(x_{g}y_{g}+\omega^{\prime})+y_{g-1}\otimes x_{g-1}(x_{g}y_{g}+\omega^{\prime})+y_{g}\otimes x_{g-1}y_{g-1}x_{g}+x_{g}\otimes x_{g-1}y_{g-1}y_{g})$
$\displaystyle\quad+c_{3}^{g-1,g}(x_{g-1}\otimes y_{g-1}(x_{g}y_{g}+\omega^{\prime})+x_{g}\otimes x_{g-1}y_{g-1}y_{g}+y_{g-1}\otimes x_{g-1}x_{g}y_{g}+y_{g}\otimes x_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime}))$
$\displaystyle\quad+c_{4}^{g-1,g}(x_{g-1}\otimes x_{g}y_{g}y_{g-1}+x_{g}\otimes y_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime})+y_{g}\otimes x_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime})+y_{g-1}\otimes x_{g-1}x_{g}y_{g})$
$\displaystyle\quad+c_{5}^{g-1,g}(x_{g-1}\otimes y_{g}y_{g-1}x_{g}+y_{g}\otimes x_{g-1}y_{g-1}x_{g}+y_{g-1}\otimes x_{g-1}(x_{g}y_{g}+\omega^{\prime})+x_{g}\otimes y_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime}))$
$\displaystyle\quad+c_{6}^{g-1,g}(x_{g-1}\otimes(x_{g}y_{g}+\omega^{\prime})y_{g-1}+y_{g}\otimes x_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime})$
$\displaystyle\qquad\qquad\quad+x_{g}\otimes y_{g}(x_{g-1}y_{g-1}+\omega^{\prime\prime})+y_{g-1}\otimes x_{g-1}(x_{g}y_{g}+\omega^{\prime}))|.$
Now, we compare the coefficients. We obtain
$\displaystyle x_{i}\otimes x_{i}y_{i}^{2}$ $\displaystyle:2c_{1}^{i}+2c_{2}^{i}=0\,(1\leq i\leq g),$ (5)
$\displaystyle x_{i}\otimes y_{i}x_{j}y_{j}$ $\displaystyle:(c_{1}^{ij}+c_{4}^{ij}+c_{5}^{ij})+(c_{2}^{ig}+c_{3}^{ig}+c_{6}^{ig})=0\,(1\leq i<j<g),$ (6)
$\displaystyle x_{i}\otimes y_{i}y_{j}x_{j}$ $\displaystyle:(c_{2}^{ij}+c_{3}^{ij}+c_{6}^{ij})-(c_{2}^{ig}+c_{3}^{ig}+c_{6}^{ig})=0\,(1\leq i<j<g),$ (7)
$\displaystyle y_{i}\otimes x_{i}x_{j}y_{j}$ $\displaystyle:(c_{1}^{ij}+c_{3}^{ij}+c_{4}^{ij})+(c_{2}^{ig}+c_{5}^{ig}+c_{6}^{ig})=0\,(1\leq i<j<g),$ (8)
$\displaystyle x_{j}\otimes x_{i}y_{i}y_{j}$ $\displaystyle:(c_{1}^{ij}+c_{2}^{ij}+c_{3}^{ij})+(c_{2}^{jg}+c_{3}^{jg}+c_{6}^{jg})=0\,(1\leq i<j<g),$ (9)
$\displaystyle x_{j}\otimes x_{i}y_{j}y_{i}$ $\displaystyle:(c_{4}^{ij}+c_{5}^{ij}+c_{6}^{ij})-(c_{2}^{jg}+c_{3}^{jg}+c_{6}^{jg})=0\,(1\leq i<j<g),$ (10)
$\displaystyle y_{j}\otimes x_{i}y_{i}x_{j}$ $\displaystyle:(c_{1}^{ij}+c_{2}^{ij}+c_{5}^{ij})+(c_{2}^{jg}+c_{5}^{jg}+c_{6}^{jg})=0\,(1\leq i<j<g),$ (11)
$\displaystyle x_{i}\otimes y_{i}x_{g}y_{g}$ $\displaystyle:c_{1}^{ig}+c_{2}^{ig}+c_{3}^{ig}+c_{4}^{ig}+c_{5}^{ig}+c_{6}^{ig}=0\,(1\leq i<g),$ (12)
$\displaystyle x_{g}\otimes x_{i}y_{i}y_{g}$ $\displaystyle:(c_{1}^{ig}+c_{2}^{ig}+c_{3}^{ig})+(c_{4}^{g-1,g}+c_{5}^{g-1,g}+c_{6}^{g-1,g})=0\,(1\leq i<g-1),$ (13)
$\displaystyle x_{g}\otimes y_{i}x_{i}y_{g}$ $\displaystyle:(c_{4}^{ig}+c_{5}^{ig}+c_{6}^{ig})-(c_{4}^{g-1,g}+c_{5}^{g-1,g}+c_{6}^{g-1,g})=0\,(1\leq i<g-1),\textrm{ and}$ (14)
$\displaystyle y_{g}\otimes x_{i}y_{i}x_{g}$ $\displaystyle:(c_{1}^{ig}+c_{2}^{ig}+c_{5}^{ig})+(c_{3}^{g-1,g}+c_{4}^{g-1,g}+c_{6}^{g-1,g})=0\,(1\leq i<g-1).$ (15)
Equations (13)-(15) holds also for $i=g-1$: (13) and (15) follows from (12), and (14) is a tautology. Now put $b^{ij}=c_{3}^{ij}-c_{5}^{ij}$. From (6) and (8), we have $b^{ij}-b^{ig}=0$ for $1\leq i<j<g$. From (9) and (11), we have $b^{ij}+b^{jg}=0$ for $1\leq i<j<g$. From (13) and (15), we have $b^{ig}-b^{g-1,g}=0$ for $1\leq i<g$. Now that we have assumed $g\geq 3$, we have
$-b^{2g}=b^{12}=b^{1g}=b^{g-1,g}=b^{2g},$
which implies $b^{2g}=0$ since $\mathrm{char}(\mathbb{K})\neq 2$. Hence we have $b^{ij}=0$ for $1\leq i<j\leq g$, which says $c_{3}^{ij}=c_{5}^{ij}$ for $1\leq i<j\leq g$. By subtracting a scalar multiple of $x_{i}y_{i}\omega$ and $y_{i}x_{i}\omega$ for $1\leq i<g$, we can freely set $c_{1}^{ig}$ and $c_{6}^{ig}$ to any value. We set $c_{1}^{ig}=c_{6}^{ig}=-c_{2}^{ig}-c_{3}^{ig}$ for $1\leq i<g$. Then, from (6)-(11), we have
$c_{1}^{ij}+c_{4}^{ij}+c_{5}^{ij}=c_{2}^{ij}+c_{3}^{ij}+c_{6}^{ij}=c_{1}^{ij}+c_{3}^{ij}+c_{4}^{ij}=c_{1}^{ij}+c_{2}^{ij}+c_{3}^{ij}=c_{4}^{ij}+c_{5}^{ij}+c_{6}^{ij}=c_{1}^{ij}+c_{2}^{ij}+c_{5}^{ij}=0.$
for $1\leq i<j<g$. This gives $c_{1}^{ij}=c_{6}^{ij}$ and $c_{2}^{ij}=c_{4}^{ij}$ for $1\leq i<j<g$. Furthermore, since we set $c_{1}^{ig}+c_{2}^{ig}+c_{3}^{ig}=0$, combining with (12), we have $c_{4}^{ig}+c_{5}^{ig}+c_{6}^{ig}=0$ for any $i$. Then, by (13) and (14), we have
$c_{1}^{ig}+c_{2}^{ig}+c_{3}^{ig}=c_{4}^{ig}+c_{5}^{ig}+c_{6}^{ig}=0$
and, together with $c_{1}^{ig}=c_{6}^{ig}$ and $c_{3}^{ig}=c_{5}^{ig}$ we already know, we have $c_{2}^{ig}=c_{4}^{ig}$ for all $i$. In conclusion, we have $c_{1}^{i}+c_{2}^{i}=0$ for all $i$ and $c_{1}^{ij}=c_{6}^{ij}$, $c_{2}^{ij}=c_{4}^{ij}$, $c_{3}^{ij}=c_{5}^{ij}$, $c_{1}^{ij}+c_{2}^{ij}+c_{3}^{ij}=0$ for $1\leq i<j\leq g$. The rest is the same as before.∎
Remark 7.5.
The author conjectures that $\operatorname{Ker}(|\bar{\Delta}|)^{(5)}=|HL^{(4)}|\oplus(\wedge^{5}H)$, whose rigorous proof could be given by a straightforward generalisation of the above procedure. The equality is already verified by a computer-assisted calculation.
We conclude this section with a conjecture.
Conjecture 7.6.
The map $\operatorname{Ker}(|\bar{\Delta}|)^{(d)}\to\operatorname{Ker}(|\bar{\Delta}_{\omega}|)^{(d)}$ induced from the natural surjection $|T(H)|\twoheadrightarrow|T(H)_{\omega}|$ is also a surjection except for $(g,d)=(2,4)$.